### v2

# Multiperiodic Processes: Ergodic Sources with a Sublinear Entropy

Multiperiodic Processes:
Ergodic Sources with a Sublinear Entropy
Łukasz Debowski Łukasz Debowski is with the Institute of Computer Science, Polish Academy of Sciences, ul. Jana Kazimierza 5, 01-248 Warszawa, Poland, e-mail: ldebowsk@ipipan.waw.pl.
We construct multiperiodic processes—a simple example of stationary ergodic stochastic processes over natural numbers that enjoy the vanishing entropy rate under a mild condition. Multiperiodic processes are supported on randomly shifted deterministic sequences called multiperiodic sequences, which can be efficiently generated using an algorithm called the Infinite Clock. Under a suitable parameterization, multiperiodic sequences exhibit relative frequencies of particular numbers given by Zipf‘s law. Exactly in the same setting, the respective multiperiodic processes satisfy an asymptotic power-law growth of block entropy, called Hilberg‘s law. Hilberg‘s law is deemed to hold for statistical language models, in particular.
Key words: ergodic processes, periodic sequences, zero entropy rate, Zipf‘s law, Hilberg‘s law
1 Introduction
The aim of this paper is to construct a simple stationary ergodic source with the vanishing entropy rate. In particular, we are interested in the block entropy obeying an asymptotic power law growth.
The notations are as follows. Let $H(X):=\operatorname{\mathbf{E}}\left[-\log P(X)\right]$ be the Shannon entropy of a discrete random variable $X$ where $\log$ is the base 2 logarithm, let $H(X|Y):=H(X,Y)-H(Y)$ be the conditional entropy, and let $I(X;Y):=H(X)-H(X|Y)$ be the mutual information. We assume a general familiarity with these Shannon information measures [1]. Relation $\sim$ denotes the equality of the asymptotic power law growth,
$f(n)\sim g(n)\stackrel{{\scriptstyle\text{def}}}{{\iff}}\lim_{n\to\infty}\frac{\log f(n)}{\log g(n)}=1.$ (1)
Blocks of random variables are denoted $X_{j}^{k}:=(X_{j},X_{j+1},\ldots,X_{k})$.
We are seeking for mathematical toy examples of stationary processes $(X_{t})_{t\in\mathbb{Z}}$ that obey
$H(X_{1}^{t})-ht\sim t^{\beta},$ (2)
where $\beta\in(0,1)$ and $h=\lim_{t\to\infty}H(X_{1}^{t})/t$ is the Shannon entropy rate. The rough proportionality (2) is called Hilberg‘s law after [2] and is deemed to hold for natural language [2, 3, 4, 5]. As it will be detailed in a separate paper, Hilberg‘s law implies the more famous neural scaling law, which is a power-law decay of the cross entropy rate of large language models [6, 7, 8].
Some simple known examples of processes that obey Hilberg‘s law (2) are Santa Fe processes [9, 4, 10]. The Santa Fe processes come in a number of varieties. Their common idea is to decompose each token $X_{t}$ as a pair of a natural number $K_{t}$ and an additional bit—which is copied from a certain sequence of bits $(Z_{k})_{k\in\mathbb{N}}$ by taking the bit at position $K_{t}$. In other words, each text token $X_{t}$ may be written as a pair
$X_{t}=(K_{t},Z_{K_{t}}),$ (3)
where $(Z_{k})_{k\in\mathbb{N}}$ is the sequence of bits and $(K_{t})_{t\in\mathbb{Z}}$ is the sequence of numbers.
The Santa Fe decomposition (3) is interesting for its abstract semantic interpretation, see [4, 5]. We may imagine that pairs $(K_{t},Z_{K_{t}})$ are logical statements that describe a certain unknown binary sequence $(Z_{k})_{k\in\mathbb{N}}$ at random but in a non-contradictory way. Namely, if statements $(k,Z_{k})$ and $(k^{\prime},Z_{k^{\prime}})$ describe the same bit $(k=k^{\prime})$ then they assert the same value $(Z_{k}=Z_{k^{\prime}})$. Thus, we may imagine that sequence $(Z_{k})_{k\in\mathbb{N}}$ expresses some unbounded immutable general knowledge which is only partially accessed and described in finite texts. For this reason, bits $Z_{k}$ appearing in representation (3) were called the ’’(elementary) facts‘‘, whereas sequence $(K_{t})_{t\in\mathbb{Z}}$ was called the ’’narration (process)‘‘ [5]. Thus, we may say that formula (3) decomposes a corpus of texts $(X_{t})_{t\in\mathbb{Z}}$ into ’’knowledge‘‘ $(Z_{k})_{k\in\mathbb{N}}$ and ’’narration‘‘ $(K_{t})_{t\in\mathbb{Z}}$.
A few previous discussions of Santa Fe processes assumed that narration $(K_{t})_{t\in\mathbb{Z}}$ is a memoryless source with the marginal Zipf-Mandelbrot distribution
$P(K_{t}=k)\sim k^{-\alpha},$ (4)
where $\alpha>1$ [11, 12]. Moreover, it was assumed that knowledge $(Z_{k})_{k\in\mathbb{N}}$ is a sequence of independent fair coin flips, independent of process $(K_{t})_{t\in\mathbb{Z}}$. Under these conditions, we obtain Hilberg‘s law (2) with $\beta=1/\alpha$. The reason for this law is as follows. Denote the random set of types [^1] We obey the distinction of terms “type” and “token” borrowed from quantitative linguistics [13, 14, 15]: A token is an element of a sequence with possible repetitions, whereas a type is an element of a set, which excludes repetitions.
$\mathcal{V}_{t}:=\left\{K_{1},K_{2},\ldots,K_{t}\right\}.$ (5)
For the Zipf-Mandelbrot distribution (4), we have the Herdan-Heaps law $\operatorname{\mathbf{E}}\operatorname{\#}\mathcal{V}_{t}\sim t^{1/\alpha}$ [13, 16, 15]. Consequently, we may express
$H(X_{1}^{t})-H(K_{1}^{t})=H(\left\{(k,Z_{k}):k\in\mathcal{V}_{t}\right\}|K_{1}^{t})=\operatorname{\mathbf{E}}\operatorname{\#}\mathcal{V}_{t}\sim t^{1/\alpha},$ (6)
which implies $H(K_{1}^{t})=H(K_{0})t=ht$, thus recovering Hilberg‘s law (2).
The above described Santa Fe process is non-ergodic and takes values in a countably infinite alphabet. The reason for its non-ergodicity is that coin flips $(Z_{k})_{k\in\mathbb{N}}$ are measurable with respect to the shift-invariant algebra of process $(X_{t})_{t\in\mathbb{Z}}$. Thus each particular value of sequence $(Z_{k})_{k\in\mathbb{N}}$ defines a distinct ergodic component [9, 5]. However, replacing individual coin flips $Z_{k}$ with time-dependent but slowly evolving Markov chains $(Z_{kt})_{t\in\mathbb{Z}}$, we obtain processes which are (strongly) mixing hence ergodic [17]. Coding natural numbers as binary strings, we obtain processes over a finite alphabet [4, 18]. All these examples enjoy Hilberg‘s law (2) with a positive rate $h>0$.
To complete the collection of invented sources, we have been interested in finding a similarly simple example of a process with Hilberg‘s law (2) and the vanishing Shannon entropy rate, $h=0$. In [19, 5], we have constructed the random hierarchical association processes (RHA), which satisfy (2) with $h=0$ but are probably non-ergodic. The construction of RHA processes is so complicated that we do not reproduce it here. In the end, we desired a simpler construction that would be additionally ergodic.
Our curiosity in such asymptotically deterministic models was initially driven by the observation that the maximal repetition length for natural language obeys a cube logarithmic growth [20]. Consequently, the conditional Rényi entropy rate may vanish for natural language [21]. Later, our interest was reinforced by the research in the neural scaling law [6, 7, 8], which suggested that Hilberg‘s law (2) holds for large language models with $h=0$. We are agnostic about the actual value of particular entropy rates for natural language but we have wished to better understand the mathematical construction of zero entropy rate processes.
This paper supplies another toy example called multiperiodic processes. Multiperiodic processes $(K_{t})_{t\in\mathbb{Z}}$ are asymptotically deterministic sequences of random natural numbers. They are not mixing but they are ergodic and, under a certain parameterization, they satisfy Hilberg‘s law with the vanishing Shannon entropy rate, namely,
$H(K_{1}^{t})\sim t^{\beta}.$ (7)
The essential idea of these processes is to define a probability measure on a set of efficiently computable but randomly shifted deterministic sequences, called multiperiodic sequences. Under a certain parameterization, the relative frequencies of particular numbers in these sequences approach the Zipf-Mandelbrot distribution (4). This suffices to imply Hilberg‘s law (7) with exponent $\beta=1/\alpha$ by a reasoning similar to bound (6). Namely, for the set of types (5), we still have the Herdan-Heaps law $\operatorname{\mathbf{E}}\operatorname{\#}\mathcal{V}_{t}\sim t^{1/\alpha}$ and we may express
$H(K_{1}^{t})=H(\left\{(k,\Sigma_{k}):k\in\mathcal{V}_{t}\right\})\sim\operatorname{\mathbf{E}}\operatorname{\#}\mathcal{V}_{t}\sim t^{1/\alpha},$ (8)
where $(\Sigma_{k})_{k\in\mathbb{N}}$ are certain random variables called seeds. Seeds parameterize random shifts of the underlying deterministic multiperiodic sequences. Variables $(\Sigma_{k})_{k\in\mathbb{N}}$ are probabilistically independent but are not measurable with respect to the shift-invariant algebra of process $(K_{t})_{t\in\mathbb{Z}}$. Thus the multiperiodic process $(K_{t})_{t\in\mathbb{Z}}$ can be ergodic.
The organization of the paper is as follows. In Section 2, we define two objects: multiperiodic sequences in Section 2.1 and multiperiodic processes in Section 2.2. We analyze their statistics in Section 3: the relative frequencies in Section 3.1, the waiting times in Section 3.2, and the number of types in Section 3.3. In Section 4, we inspect the information flow: estimation of seeds given a finite sample in Section 4.1 and the block entropy in Section 4.2. These general developments are supplemented by examples in Section 5: constant periods in Section 5.1 and linear periods in Section 5.2. Section 6 concludes the article.
2 Definitions
2.1 Multiperiodic sequences
Multiperiodic sequences are an idea generalized from the work by Kalociński and Steifer [22].
Definition 1 (multiperiodic sequence)
Consider a sequence $(\pi_{k})_{k\in\mathbb{N}}$ of natural numbers $\pi_{k}\geq 2$. Let $(\sigma_{k})_{k\in\mathbb{N}}$ be another sequence of natural numbers $\sigma_{k}\in\left\{1,2,\ldots,\pi_{k}\right\}$ such that set $\left\{k\in\mathbb{N}:\sigma_{k}=1\right\}$ is infinite. Parameters $\pi_{k}$ are called periods, whereas parameters $\sigma_{k}$ are called seeds. Given $(\pi_{k})_{k\in\mathbb{N}}$ and $(\sigma_{k})_{k\in\mathbb{N}}$, the multiperiodic sequence $(k_{t})_{t\in\mathbb{Z}}$ is the result of the initial condition
$\phi_{k,1}=\sigma_{k}$ (9)
and recursion
$k_{t}$ $=\min\left\{k\in\mathbb{N}:\phi_{k,t}=1\right\},$ (10)
$\phi_{k,t+1}$ =$=\begin{cases}\phi_{k,t}-1,&k<k_{t},\\ \pi_{k},&k=k_{t},\\ \phi_{k,t},&k>k_{t},\end{cases}$ (11)
which can be inverted as
$k_{t}$ $=\min\left\{k\in\mathbb{N}:\phi_{k,t+1}=\pi_{k}\right\},$ (12)
$\phi_{k,t}$ =$=\begin{cases}\phi_{k,t+1}+1,&k<k_{t},\\ 1,&k=k_{t},\\ \phi_{k,t+1},&k>k_{t}.\end{cases}$ (13)
The auxiliary matrix $(\phi_{k,t})_{k\in\mathbb{N},t\in\mathbb{Z}}$ is called the hand matrix.
Here we present two examples of multiperiodic sequences.
Example 1 (seemingly A001511 in [23])
For a solution of a different problem in prediction, Kalociński and Steifer [22] considered a particular multiperiodic sequence $(k_{t})_{t\in\mathbb{N}}$ with constant periods $\pi_{k}=2$ and seeds $\sigma_{k}=1$, namely,
$1,2,1,3,1,2,1,4,1,2,1,3,1,2,1,5,1,2,1,3,1,2,1,4,1,2,1,\ldots$ (14)
Our applications are closer to this example:
Example 2 (seemingly A028920 in [23])
Let $\pi_{k}=1+k$ and $\sigma_{k}=1$ for all $k\in\mathbb{N}$. Then the multiperiodic sequence $(k_{t})_{t\in\mathbb{N}}$ is
$1,2,1,3,1,4,1,2,1,5,1,6,1,2,1,3,1,7,1,2,1,8,1,4,1,2,1,\ldots$ (15)
As we show in Section 5.2, for periods $\pi_{k}\approx ck$, the relative frequency of number $k$ approaches the Zipf-Mandelbrot law (4) with $\alpha=(c+1)/c$.
Recursion (9)–(13) for $k\in\mathbb{N}$ can be rewritten as Algorithm 1, called the Infinite Clock. The Infinite Clock operates with infinite data structures for simplicity of presentation. It works like a clock with infinitely many hands—indexed by natural numbers—that move counterclockwise, hence the name. As in the usual clock, each hand moves with a different speed, dictated by its specific period. The initial positions of hands are given by seeds. The algorithm outputs the value of a hand index whenever the hand passes noon.
Algorithm 1 The Infinite Clock
1:Sequence of $\pi_{k}\in\mathbb{N}$ for $k\in\mathbb{N}$. $\triangleright$ The periods.
2:Sequence of $\phi_{k}:=\sigma_{k}\in\left\{1,2,\ldots,\pi_{k}\right\}$ for $k\in\mathbb{N}$. $\triangleright$ The hands.
3:Sequence of $k_{t}\in\mathbb{N}$ for $t\in\mathbb{N}$. $\triangleright$ The multiperiodic sequence.
4:for $t\in\mathbb{N}$ do
5:     $k^{*}:=0$
6:     $k:=1$
7:     while $k^{*}=0$ do
8:         if $\phi_{k}>1$ then
9:              $\phi_{k}:=\phi_{k}-1$
10:         else
11:              $k^{*}:=k$
12:         $k:=k+1$
13:     $\phi_{k^{*}}:=\pi_{k^{*}}$
14:     $k_{t}:=k^{*}$
To be precise, Algorithm 1 operates with hand variables $(\phi_{k})_{k\in\mathbb{N}}$, where each hand $\phi_{k}$ takes values in set $\left\{1,2,\ldots,\pi_{k}\right\}$ and is initialized with the value of the corresponding seed, $\phi_{k}:=\sigma_{k}$. To determine the value of the next element $k_{t}$ of the sequence, the algorithm decrements consecutive hands $(\phi_{k})_{k\in\mathbb{N}}$ starting from index $k=1$ until it finds an index $k^{*}$ such that $\phi_{k^{*}}=1$. Then the loop is broken, this particular hand is reset to its own maximal value $\phi_{k^{*}}:=\pi_{k^{*}}$, and the corresponding index $k^{*}$ of the hand is output as the next element of the sequence $k_{t}:=k^{*}$.
Let us observe that the computation falls into an infinite loop in lines 4–9 of Algorithm 1 if and only if there is no $k\in\mathbb{N}$ such that $\phi_{k}=1$. This happens precisely if and only if set $\left\{k\in\mathbb{N}:\sigma_{k}=1\right\}$ is finite, which is excluded by Definition 1. The easiest way to prevent this situation is to set the minimal seeds $\sigma_{k}=1$ for all $k\in\mathbb{N}$.
As a general fact, we notice a time-reversal symmetry.
Theorem 1
For the multiperiodic sequence $(k_{t})_{t\in\mathbb{Z}}$ with periods $\pi_{k}$ and seeds $\sigma_{k}$, sequence $(k_{1-t})_{t\in\mathbb{Z}}$ is the multiperiodic sequence with periods $(\pi_{k})_{k\in\mathbb{N}}$ and seeds $(\sigma^{R}_{k})_{k\in\mathbb{N}}$ defined as $\sigma^{R}_{k}:=\pi_{k}-\sigma_{k}+1$.
  • Proof: Let $(\phi^{R}_{k,t})_{k\in\mathbb{N},t\in\mathbb{Z}}$ and $(k^{R}_{t})_{t\in\mathbb{Z}}$ be defined analogously as (10)–(13) but with the initial condition $\phi^{R}_{k,1}=\sigma^{R}_{k}$. Then by induction we obtain $\phi^{R}_{k,t}=\pi_{k}-\phi_{k,1-t}+1$ and $k^{R}_{t}=k_{1-t}$. $\Box$
Moreover, we observe a partial periodicity of decimated multiperiodic sequences, which are also multiperiodic.
Theorem 2
Let $(k^{\geq r}_{t})_{t\in\mathbb{N}}$ be the subsequence of the multiperiodic sequence $(k_{t})_{t\in\mathbb{N}}$ with periods $\pi_{k}$ and seeds $\sigma_{k}$ from which we have removed all tokens $k_{t}<r$. We claim that:
  1. 1.
    Sequence $(k^{\geq r}_{t}-r+1)_{t\in\mathbb{N}}$ is the multiperiodic sequence with periods $\pi^{\geq r}_{k}=\pi_{k-r+1}$ and seeds $\sigma^{\geq r}_{k}=\sigma_{k-r+1}$.
  2. 2.
    We have $k^{\geq r}_{t}=r\iff t\equiv\sigma_{r}\mod\pi_{r}$.
  • Proof: We proceed by induction on $r$. As for the initial step, we have $(k^{\geq 1}_{t})_{t\in\mathbb{N}}=(k_{t})_{t\in\mathbb{N}}$. Thus claim 1. holds by definition, whereas claim 2. follows since $k_{t}=1$ if and only if $t\equiv\sigma_{1}\mod\pi_{1}$. As for the inductive step, it is sufficient to show that claim 1. holds, whereas claim 2. follows by claim 1. for the inductive step combined with claim 2. for the initial step. The proof of claim 1. for the inductive step rests on the fact that to obtain $(k^{\geq r}_{t})_{t\in\mathbb{N}}$ instead of $(k_{t})_{t\in\mathbb{N}}$, it suffices to initialize $k:=r$ instead of $k:=1$ in line 3 of Algorithm 1. $\Box$
2.2 Multiperiodic processes
Multiperiodic sequences can be applied to define stationary ergodic processes. The idea is to apply a random initialization, that is, uniformly random seeds. This leads to this construction, called a multiperiodic process.
Definition 2 (multiperiodic process)
The multiperiodic process $(K_{t})_{t\in\mathbb{Z}}$ with periods $(\pi_{k})_{k\in\mathbb{N}}$ is the random multiperiodic sequence with periods $(\pi_{k})_{k\in\mathbb{N}}$ and probabilistically independent random seeds $(\Sigma_{k})_{k\in\mathbb{N}}$ with uniform distributions
$P(\Sigma_{k}=\sigma_{k})=\frac{1}{\pi_{k}},\quad\sigma_{k}\in\left\{1,2,\ldots,\pi_{k}\right\}$ (16)
provided set $\left\{k\in\mathbb{N}:\Sigma_{k}=1\right\}$ is infinite almost surely.
We have a simple criterion of existence of a multiperiodic process.
Theorem 3
The multiperiodic process exists if and only if its periods satisfy
$\sum_{k=1}^{\infty}\frac{1}{\pi_{k}}=\infty.$ (17)
  • Proof: Observe that events $(\Sigma_{k}=1)$ are independent. Hence by the Borel-Cantelli lemmas [24, Theorems 4.3 and 4.4], set $\left\{k\in\mathbb{N}:\Sigma_{k}=1\right\}$ is infinite almost surely if and only if
    $\sum_{k=1}^{\infty}P(\Sigma_{k}=1)=\infty.$ (18)
    In view of the uniform distributions of $(\Sigma_{k})_{k\in\mathbb{N}}$, we obtain the claim. $\Box$
Moreover, each multiperiodic process admits a frequency interpretation of its probability distribution. Namely, it is stationary and ergodic.
Theorem 4
The multiperiodic process $(K_{t})_{t\in\mathbb{Z}}$ is stationary and ergodic but is not mixing.
  • Proof: A sequence $(a_{t})_{t\in\mathbb{Z}}$ is called $p$-periodic if
    $(a_{t+s})_{t\in\mathbb{Z}}=(a_{t})_{t\in\mathbb{Z}}\iff s\equiv 0\mod p.$ (19)
    A stochastic process $(A_{t})_{t\in\mathbb{Z}}$ is called $p$-periodic if
    $P((A_{t})_{t\in\mathbb{Z}}=(a_{t+s})_{t\in\mathbb{Z}})=\frac{1}{p},\quad s\geq 0,$ (20)
    for a certain $p$-periodic sequence $(a_{t})_{t\in\mathbb{Z}}$. It is known that $p$-periodic processes are stationary and ergodic but they are not mixing [5, Section 4.1].
    Let an $m\in\mathbb{N}$. Let $K^{\leq m}_{t}:=\min\left\{K_{t},m\right\}$. Observe that process $(K^{\leq m}_{t})_{t\in\mathbb{Z}}$ depends only on seeds $(\Sigma_{k})_{k\leq m}$. Denote $p_{m}:=\prod_{k=1}^{m}\pi_{k}$. We will show that process $(K^{\leq m}_{t})_{t\in\mathbb{Z}}$ is $p$-periodic where $p\leq p_{m}$.
    Let $(\Phi_{k,t})_{k\in\mathbb{N},t\in\mathbb{Z}}$ be the hand matrix of sequence $(K_{t})_{t\in\mathbb{Z}}$. By recursion (10)–(13), there is a bijection $(\Phi_{k,t})_{k\leq m}\leftrightarrow(\Phi_{k,t+1})_{k\leq m}$. Hence process $(\Phi_{k,t})_{k\leq m,t\in\mathbb{Z}}$ is a Markov process over time $t$. There are exactly $p_{m}$ distinct states of hands $(\phi_{k})_{k\leq m}$ and it can be shown that all states communicate. Hence Markov process $(\Phi_{k,t})_{k\leq m,t\in\mathbb{Z}}$ is irreducible, each state $(\phi_{k})_{k\leq m}$ is visited exactly once within the period $p_{m}$, and the uniform marginal distribution of the initial state $(\Phi_{k,1})_{k\leq m}=(\Sigma_{k})_{k\leq m}$ is the invariant distribution. In turn, process $(\Phi_{k,t})_{k\leq m,t\in\mathbb{Z}}$ is stationary, ergodic, and $p_{m}$-periodic. Since each $K^{\leq m}_{t}$ is a function of $(\Phi_{k,t})_{k\leq m}$, hence process $(K^{\leq m}_{t})_{t\in\mathbb{Z}}$ is $p$-periodic where $p\leq p_{m}$.
    Thus process $(K^{\leq m}_{t})_{t\in\mathbb{Z}}$ is stationary and ergodic but is not mixing. Since $m\in\mathbb{N}$ is arbitrarily chosen, hence also process $(K_{t})_{t\in\mathbb{Z}}$ is stationary and ergodic but is not mixing. The reason is that verifying stationarity, ergodicity, and mixing can be reduced to checking probabilities of finite blocks by [5, Theorem 4.10] or [25, Lemma 7.15]. $\Box$
3 Statistics
3.1 Relative frequencies
In the following, we analyze several statistics of multiperiodic sequences and multiperiodic processes. We begin with the relative frequencies and the marginal distribution.
Theorem 5
Let $(k_{t})_{t\in\mathbb{Z}}$ be a multiperiodic sequence with periods $(\pi_{k})_{k\in\mathbb{N}}$. The relative frequencies exist and equal
$\lim_{t\to\infty}\frac{1}{t}\sum_{j=1}^{t}{\bf 1}{\left\{k_{j}\geq k\right\}}=f_{k}:=\prod_{j=1}^{k-1}\frac{\pi_{j}-1}{\pi_{j}}.$ (21)
  • Proof: Let $(k_{t}^{\geq r})_{t\in\mathbb{N}}$ be the subsequence of $(k_{t})_{t\in\mathbb{N}}$ from which we have removed all tokens $k_{t}<r$. We proceed by induction on $k$. First, we observe
    $\lim_{t\to\infty}\frac{1}{t}\sum_{j=1}^{t}{\bf 1}{\left\{k_{j}\geq 1\right\}}=1=f_{1}.$ (22)
    Second, in view of claim 2. of Theorem 2, we notice
    $\lim_{t\to\infty}\frac{1}{t}\sum_{j=1}^{t}{\bf 1}{\left\{k_{j}\geq k\right\}}$
    $=\left(\lim_{t\to\infty}\frac{1}{t}\sum_{j=1}^{t}{\bf 1}{\left\{k_{j}\geq k-1\right\}}\right)\left(\lim_{t\to\infty}\frac{1}{t}\sum_{j=1}^{t}{\bf 1}{\left\{k_{j}^{(k-1)}\geq k\right\}}\right)$
    $=f_{k-1}\left(\frac{\pi_{k-1}-1}{\pi_{k-1}}\right)=f_{k}.$ (23)
    Hence the claim follows. $\Box$
Theorem 6
The multiperiodic process $(K_{t})_{t\in\mathbb{Z}}$ with periods $(\pi_{k})_{k\in\mathbb{N}}$ has the marginal distribution
$P(K_{0}=k)=p_{k}:=f_{k}-f_{k+1},$ (24)
where $f_{k}$ is the marginal distribution function (21).
  • Proof: We obtain distribution (24) by the Birkhoff ergodic theorem, which establishes the almost sure equality of probabilities and relative frequencies for stationary ergodic processes [26, 27]. $\Box$
3.2 Waiting times
Another important statistic of a stochastic process is the waiting time. We define the waiting time as
$W_{k}:=\inf\left\{t\in\mathbb{N}:K_{t}=k\right\}.$ (25)
The Kac lemma [28] states that
$\operatorname{\mathbf{E}}\left(W_{k}|K_{0}=k\right)=\frac{1}{P(K_{0}=k)}$ (26)
for a stationary ergodic process $(K_{t})_{t\in\mathbb{Z}}$. However, for a multiperiodic process, we can obtain a stronger uniform bound.
Theorem 7
For a multiperiodic process $(K_{t})_{t\in\mathbb{Z}}$ with periods $(\pi_{k})_{k\in\mathbb{N}}$, we have a uniform bound
$W_{k}\leq w_{k}:=\sum_{j=2}^{k}\frac{1}{f_{j}}+\frac{\pi_{k}}{f_{k}},$ (27)
where $f_{k}$ is the marginal distribution function (21).
  • Proof: Consider a prefix $(K_{1},K_{2},\ldots,K_{T_{1}})$ of the multiperiodic process. When we delete all $K_{t}<r$, the sequence shortens to $(K^{\geq r}_{1},K^{\geq r}_{2},\ldots,K^{\geq r}_{T_{r}})$. We observe that $W_{k}\leq T_{1}$ if $T_{k}=\pi_{k}$, whereas the partial periodicity implies
    $T_{r}-T_{r+1}\leq\frac{T_{r}}{\pi_{r}}+1.$ (28)
    Hence we have
    $T_{r}\leq(T_{r+1}+1)\cdot\frac{\pi_{r}}{\pi_{r}-1}$ (29)
    and consequently we obtain
    $W_{k}$ $\leq\left(\ldots\left(\left(\left(\pi_{k}+1\right)\frac{\pi_{k-1}}{\pi_{k-1}-1}+1\right)\frac{\pi_{k-2}}{\pi_{k-2}-1}+1\right)\ldots\right)\frac{\pi_{1}}{\pi_{1}-1}$
    $=\sum_{j=2}^{k}\frac{1}{f_{j}}+\frac{\pi_{k}}{f_{k}}=w_{k}$ (30)
    in view of identity
    $\left(\ldots\left(\left(\left(a_{k}+1\right)a_{k-1}+1\right)a_{k-2}+1\right)\ldots\right)a_{1}=\sum_{j=1}^{k}\prod_{i=1}^{j}a_{i}.$ (31)
    $\Box$
By the way, we notice this fact.
Definition 3 (increasing periods)
Periods $(\pi_{k})_{k\in\mathbb{N}}$ are called increasing if $\pi_{k+1}\geq\pi_{k}$ for all $k\in\mathbb{N}$.
Theorem 8
For increasing periods $(\pi_{k})_{k\in\mathbb{N}}$, we have:
$w_{k+1}$ $\geq w_{k},$ (32)
$P(K_{0}=k+1)$ $\leq P(K_{0}=k).$ (33)
  • Proof: We have $w_{k+1}\geq w_{k}$ since
    $\pi_{k}+1\leq\pi_{k+1}+1\leq(\pi_{k+1}+1)\frac{\pi_{k}}{\pi_{k}-1}+1.$ (34)
    Similarly $P(K_{0}=k)=f_{k}-f_{k+1}\geq f_{k+1}-f_{k+2}=P(K_{0}=k+1)$ since
    $1-\frac{\pi_{k}-1}{\pi_{k}}\geq\frac{\pi_{k}-1}{\pi_{k}}\left(1-\frac{\pi_{k+1}-1}{\pi_{k+1}}\right).$ (35)
    $\Box$
3.3 Number of types
Let us denote the random set of observed types
$\mathcal{V}_{t}:=\left\{K_{1},K_{2},\ldots,K_{t}\right\}=\left\{k\in\mathbb{N}:W_{k}\leq t\right\}.$ (36)
We denote its cardinality
$V_{t}:=\operatorname{\#}\mathcal{V}_{t}.$ (37)
Let us introduce also three more statistics:
$u_{t}$ $:=\min\left\{k\in\mathbb{N}:w_{k}>t\right\},$ (38)
$U_{t}$ $:=\min\left\{k\in\mathbb{N}:k\not\in\mathcal{V}_{t}\right\},$ (39)
$g_{t}$ $:=\max\left\{k\in\mathbb{N}:P(K_{0}=k)\geq\frac{1}{t}\right\},$ (40)
$q_{t}$ $:=P(K_{0}>g_{t}).$ (41)
Theorem 9
For a multiperiodic process $(K_{t})_{t\in\mathbb{Z}}$, we have
$u_{t}-1\leq U_{t}-1\leq V_{t}.$ (42)
  • Proof: The inequalities follow by $w_{k}\geq W_{k}$ and $U_{t}-1\leq\operatorname{\#}\mathcal{V}_{t}$. $\Box$
Applying the idea by [14] for memoryless sources, we can prove this proposition.
Theorem 10
For a stationary process $(K_{t})_{t\in\mathbb{Z}}$ over natural numbers, we have
$\operatorname{\mathbf{E}}V_{t}\leq g_{t}+tq_{t}.$ (43)
  • Proof: We observe
    $\operatorname{\mathbf{E}}V_{t}=\operatorname{\mathbf{E}}\sum_{k=1}^{\infty}{\bf 1}{\left\{k\in\mathcal{V}_{t}\right\}}=\sum_{k=1}^{\infty}P(k\in\mathcal{V}_{t})\leq g_{t}+\sum_{k>g_{t}}P(k\in\mathcal{V}_{t})$ (44)
    and the union bound for stationary processes
    $P(k\in\mathcal{V}_{t})=P(K_{1}=k\lor\ldots\lor K_{t}=k)\leq tP(K_{0}=k).$ (45)
    Hence the claim follows. $\Box$
Quotient $V_{t}/t$ is called the type-token ratio in quantitative linguistics [13, 14, 15]. It is easy to see that it tends to zero in the stationary case.
Theorem 11
For a stationary process $(K_{t})_{t\in\mathbb{Z}}$ over natural numbers,
$\lim_{t\to\infty}\frac{\operatorname{\mathbf{E}}V_{t}}{t}$ $=0.$ (46)
  • Proof: We observe that bound
    $\operatorname{\mathbf{E}}V_{t}\leq h_{t}+tP(K_{0}>h_{t})$ (47)
    holds for any function $h_{t}$. In particular, for an $\epsilon>0$, we may take $h_{t}=\epsilon t/2$. For all sufficiently large $t$, we observe $P(K_{0}>h_{t})\leq\epsilon/2$. Hence, for these $t$, we have $\operatorname{\mathbf{E}}V_{t}/t\leq h_{t}/t+P(K_{0}>h_{t})\leq\epsilon$. By arbitrariness of $\epsilon$, we derive (46). $\Box$
We need a bit stronger bound.
Theorem 12
For a stationary process $(K_{t})_{t\in\mathbb{Z}}$ over natural numbers such that $\operatorname{\mathbf{E}}(\log K_{0})^{3}<\infty$, we have
$\lim_{t\to\infty}\frac{\operatorname{\mathbf{E}}V_{t}\log t}{t}$ $=0,$ (48)
[…]
  • Proof: Observe by the Markov inequality that
    $\sum_{k=1}^{\infty}kP\left(K_{0}\geq\frac{2^{k}}{k^{3}}\right)$ $\leq\sum_{k=1}^{\infty}kP\left((\log K_{0})^{3}\geq(k-3\log k)^{3}\right)$
    $\leq\operatorname{\mathbf{E}}(\log K_{0})^{3}\sum_{k=1}^{\infty}\frac{k}{(k-3\log k)^{3}}<\infty.$ (50)
    Choose $h_{t}=t/(\log t)^{3}$ in (47). Then we obtain
    $\frac{\operatorname{\mathbf{E}}V_{2^{k}}k}{2^{k}}\leq\frac{1}{k^{2}}+kP\left(K_{0}\geq\frac{2^{k}}{k^{3}}\right).$ (51)
    Consequently, by the Markov inequality, we have
    $\sum_{k=1}^{\infty}P\left(\frac{V_{2^{k}}k}{2^{k}}\geq\epsilon\right)$ $\leq\frac{1}{\epsilon}\sum_{k=1}^{\infty}\frac{\operatorname{\mathbf{E}}V_{2^{k}}k}{2^{k}}$
    $\leq\frac{1}{\epsilon}\sum_{k=1}^{\infty}\left(\frac{1}{k^{2}}+kP\left(K_{0}\geq\frac{2^{k}}{k^{3}}\right)\right)<\infty.$ (52)
    Hence $\lim_{k\to\infty}\operatorname{\mathbf{E}}V_{2^{k}}k/2^{k}=0$, whereas by the Borel-Cantelli lemma [24, Theorems 4.3], we derive $V_{2^{k}}k/2^{k}<\epsilon$ for all but finitely many $k$ almost surely. Notice now that $V_{t}$ is a growing function of $t$. Hence $(V_{t}\log t)/t\leq 2V_{2^{k}}k/2^{k}$ for $2^{k-1}\leq t\leq 2^{k}$. In particular, we have (48) and $\limsup_{t\to\infty}(V_{t}\log t)/t\leq 2\epsilon$ almost surely. By arbitrariness of $\epsilon$, we derive (49). $\Box$
4 Information
4.1 Seed estimator
Let us analyze the relationship between a finite multiperiodic sample and the pool of random seeds in more detail. For bounding the entropy of the sample, it pays off to state this proposition explicitly.
Theorem 13
Consider a multiperiodic process $(K_{t})_{t\in\mathbb{Z}}$. We claim that:
  1. 1.
    Set $\left\{(k,\Sigma_{k}):k\in\mathcal{V}_{t}\right\}$ is a function of sample $K_{1}^{t}$.
  2. 2.
    Define the seed estimator
    $\hat{\Sigma}_{k}:=\begin{cases}\Sigma_{k},&k\in\mathcal{V}_{t},\\ \pi_{k},&k\not\in\mathcal{V}_{t}.\end{cases}$ (53)
    Let $(\hat{K}_{t})_{t\in\mathbb{Z}}$ be the multiperiodic sequence with seeds $(\hat{\Sigma}_{k})_{k\in\mathbb{N}}$. Then
    $K_{1}^{t}=\hat{K}_{1}^{t}.$ (54)
  3. 3.
    Sample $K_{1}^{t}$ is a function of set $\left\{(k,\Sigma_{k}):k\in\mathcal{V}_{t}\right\}$.
  4. 4.
    If $k\in\mathcal{V}_{t}$ then $k\leq M_{t}$ and $\Sigma_{k}\leq t$, where
    $M_{t}:=\max\mathcal{V}_{t}.$ (55)
  • Proof:
    • Let $L_{k}=\min\left\{t\geq 1:K_{t}^{(k)}=k\right\}$, where $(K^{(k)}_{t})_{t\in\mathbb{N}}$ is the subsequence of sequence $(K_{t})_{t\in\mathbb{N}}$ from which we have removed all tokens $k_{t}<k$. If $k\in\mathcal{V}_{t}$ then the first $L_{k}$ tokens of $(K^{(k)}_{t})_{t\in\mathbb{N}}$ are a function of sample $K_{1}^{t}$. By claim 2. of Theorem 2, $L_{k}=\Sigma_{k}$. Hence we obtain claim 1.
    • We observe that altering $\Sigma_{k}$ for $k\not\in\mathcal{V}_{t}$ to any value in set
      $\left\{\Sigma_{k},\Sigma_{k}+1,\ldots,\pi_{k}\right\}$
      does not change the resulting multiperiodic sample $K_{1}^{t}$. It is so since any such change only delays the occurrence of token $k$ beyond the scope of $K_{1}^{t}$, leaving the tokens within $K_{1}^{t}$ intact. Hence claim 2. follows.
    • Obviously, claim 2. implies claim 3.
    • If $k\in\mathcal{V}_{t}$ then $k\leq\max\mathcal{V}_{t}$ and $\Sigma_{k}=L_{k}\leq t$ since the first $L_{k}$ tokens of sequence $(K^{(k)}_{t})_{t\in\mathbb{N}}$ are a subsequence of sample $K_{1}^{t}$. Hence we obtain claim 4.
    $\Box$
4.2 Block entropy
In this section, we sandwich bound the Shannon entropy in terms of the number of types. For this goal, we use a bound stemming from the Hölder and Jensen inequalities. Let $\ln$ denote the natural logarithm.
Lemma 1
Suppose that $X\geq 0$ and $Y\geq 1$. Then for any $p>1$, we have
$\operatorname{\mathbf{E}}X\ln Y\leq(\operatorname{\mathbf{E}}X^{p})^{1/p}\left(\ln\operatorname{\mathbf{E}}Y+\frac{1}{p-1}\right).$ (56)
  • Proof: Let $p>1$ and $q=p/(p-1)$. For $X,Z\geq 0$, the Hölder inequality is
    $\operatorname{\mathbf{E}}XZ\leq(\operatorname{\mathbf{E}}X^{p})^{1/p}(\operatorname{\mathbf{E}}Z^{q})^{1/q}.$ (57)
    The second derivative of function $y\mapsto(\ln y)^{q}$ is negative for $y>e^{q-1}$. Hence by the Jensen inequality (J), we obtain
    $\operatorname{\mathbf{E}}(\ln Y)^{q}\leq\operatorname{\mathbf{E}}(\ln(e^{q-1}Y))^{q}\stackrel{{\scriptstyle(J)}}{{\leq}}(\ln(e^{q-1}\operatorname{\mathbf{E}}Y))^{q}\leq\left(\ln\operatorname{\mathbf{E}}Y+\frac{1}{p-1}\right)^{q}.$ (58)
    Using $Z=\ln Y$ in the Hölder inequality and chaining it with the above chain of inequalities, we derive the claim. $\Box$
Subsequently, we bound the entropy. Here $\log$ is the base 2 logarithm.
Theorem 14
For a multiperiodic process $(K_{t})_{t\in\mathbb{Z}}$ and for any $q\geq 1$,
$u_{t}-1\leq H(K_{1}^{t})\leq\left[e\operatorname{\mathbf{E}}V_{t}+3\right]\left[(q+2)\log t+q\log\operatorname{\mathbf{E}}K_{0}^{1/q}\right]+2.$ (59)
  • Proof: By claim 1. of Theorem 13 and property $k<u_{t}\implies k\in\mathcal{V}_{t}$, the tuple of seeds $\Sigma_{1}^{u_{t}-1}$ is a function of sample $K_{1}^{t}$. Hence we obtain the lower bound
    $H(K_{1}^{t})\geq H(\left\{(k,\Sigma_{k}):k\in\mathcal{V}_{t}\right\})\geq H(\Sigma_{1}^{u_{t}-1})\geq u_{t}-1,$ (60)
    since seeds are independent and uniformly distributed over two or more values. As for the upper bound, by claims 3. and 4. of Theorem 13, we obtain
    $H(K_{1}^{t})$ $\leq H(\left\{(k,\Sigma_{k}):k\in\mathcal{V}_{t}\right\})$
    $\leq H(M_{t})+H(\left\{(k,\Sigma_{k}):k\in\mathcal{V}_{t}\right\}|M_{t})$
    $\leq 2\operatorname{\mathbf{E}}\log(M_{t}+1)+\operatorname{\mathbf{E}}(V_{t}+1)(\log M_{t}+\log t)$
    $\leq 2\operatorname{\mathbf{E}}\log M_{t}+\operatorname{\mathbf{E}}(V_{t}+1)(\log M_{t}+\log t)+2.$ (61)
    Next, we observe $1\leq M_{t}^{1/q}\leq\sum_{j=1}^{t}K_{j}^{1/q}$. Hence $\operatorname{\mathbf{E}}M_{t}^{1/q}\leq t\operatorname{\mathbf{E}}K_{0}^{1/q}$ by stationarity. Thus, by Lemma 1, for any $p>1$, we may write
    $H(K_{1}^{t})$ $\leq 2\operatorname{\mathbf{E}}q\log M_{t}^{1/q}+\operatorname{\mathbf{E}}(V_{t}+1)\left(q\log M_{t}^{1/q}+\log t\right)+2$
    $\leq[(\operatorname{\mathbf{E}}V_{t}^{p})^{1/p}+3]\left[q\log\operatorname{\mathbf{E}}M_{t}^{1/q}+\log t+\frac{\log e}{p-1}\right]+2$
    $\leq[(\operatorname{\mathbf{E}}V_{t}^{p})^{1/p}+3]\left[(q+1)\log t+q\log\operatorname{\mathbf{E}}K_{0}^{1/q}+\frac{\log e}{p-1}\right]+2.$ (62)
    Let us substitute $p=1+1/\ln t$ and apply inequality $V_{t}^{p}\leq t^{p-1}V_{t}$. Then we may rewrite
    $H(K_{1}^{t})$ $\leq\left[(t^{p-1}\operatorname{\mathbf{E}}V_{t})^{1/p}+3\right]\left[(q+1)\log t+q\log\operatorname{\mathbf{E}}K_{0}^{1/q}+\frac{\log e}{p-1}\right]+2$
    $=\left[(e\operatorname{\mathbf{E}}V_{t})^{\frac{\ln t}{\ln t+1}}+3\right]\left[(q+2)\log t+q\log\operatorname{\mathbf{E}}K_{0}^{1/q}\right]+2$
    $\leq\left[e\operatorname{\mathbf{E}}V_{t}+3\right]\left[(q+2)\log t+q\log\operatorname{\mathbf{E}}K_{0}^{1/q}\right]+2,$ (63)
    which is the claimed inequality. $\Box$
Hence we have a criterion for the vanishing Shannon entropy rate.
Theorem 15
Suppose that $\operatorname{\mathbf{E}}K_{0}^{1/q}<\infty$ for a $q\geq 1$. Then the multiperiodic process $(K_{t})_{t\in\mathbb{Z}}$ satisfies
$\lim_{t\to\infty}\frac{H(K_{1}^{t})}{t}$ $=0.$ (64)
  • Proof: The claim follows by (48) and (59) since $\operatorname{\mathbf{E}}K_{0}^{1/q}<\infty$ for a $q\geq 1$ implies $\operatorname{\mathbf{E}}(\log K_{0})^{3}<\infty$. $\Box$
5 Examples
5.1 Constant periods
In the following, we work out some concrete examples of processes. First, we analyze the statistics of the multiperiodic process with all periods equal to a fixed constant. This case leads to the number of types that grows logarithmically with the sample size.
Let $\pi_{k}=p$ for all $k\in\mathbb{N}$. Then we obtain
$f_{k}$ $=\prod_{j=1}^{k-1}\frac{\pi_{j}-1}{\pi_{j}}=\left(\frac{p-1}{p}\right)^{k-1},$ (65)
$p_{k}$ $=f_{k}-f_{k+1}=\frac{1}{p}\left(\frac{p-1}{p}\right)^{k-1},$ (66)
$w_{k}$ $=\sum_{j=2}^{k}\frac{1}{f_{j}}+\frac{\pi_{k}}{f_{k}}=2p\left(\frac{p}{p-1}\right)^{k-1}-1.$ (67)
Hence we may evaluate
$u_{t}$ $=\min\left\{k:w_{k}>t\right\}=\left\lceil\frac{\log(t+1)-\log p-1}{\log\left(\frac{p}{p-1}\right)}\right\rceil+2,$ (68)
$g_{t}$ $=\max\left\{k:p_{k}^{-1}\leq t\right\}=\left\lfloor\frac{\log(t+1)-\log p}{\log\left(\frac{p}{p-1}\right)}\right\rfloor+1,$ (69)
$\log q_{t}$ $=\log f_{g_{t}+1}=-g_{t}\log\left(\frac{p}{p-1}\right)\leq-\log(t+1)+\log(p).$ (70)
Hence bound $u_{t}-1\leq\operatorname{\mathbf{E}}V_{t}\leq g_{t}+tq_{t}$ can be rewritten as
$\left\lceil\frac{\log(t+1)-\log p-1}{\log\left(\frac{p}{p-1}\right)}\right\rceil+1\leq\operatorname{\mathbf{E}}V_{t}\leq\left\lfloor\frac{\log(t+1)-\log p}{\log\left(\frac{p}{p-1}\right)}\right\rfloor+1+p.$ (71)
In this case, the Shannon entropy rate vanishes but we do not have Hilberg‘s law (7)—by Theorem 14.
5.2 Linear periods
Second, we inspect the statistics of the multiperiodic process with periods that grow linearly. This case leads to the number of types that grows like a power law. The power-law growth of the number of types is known as the Herdan-Heaps law in quantitative linguistics [13, 16, 15].
Let $\pi_{k}=1+\left\lceil ck\right\rceil$ for a $c>0$ and all $k\in\mathbb{N}$. We may upper bound
$\prod_{j=1}^{k-1}\frac{\pi_{j}-1}{\pi_{j}}$ $=\prod_{j=1}^{k-1}\left(1-\frac{1}{1+\left\lceil cj\right\rceil}\right)\leq\exp\left(\sum_{j=1}^{k-1}\ln\left(1-\frac{1}{cj+2}\right)\right)$
$\leq\exp\left(-\sum_{j=1}^{k-1}\frac{1}{cj+2}\right)\leq\exp\left(-\int_{1}^{k}\frac{dx}{cx+2}\right)$
$=\exp\left(-\frac{1}{c}\ln\frac{ck+2}{c+2}\right)=\left(\frac{c+2}{ck+2}\right)^{1/c}.$ (72)
Similarly, we may bound conversely
$\prod_{j=1}^{k-1}\frac{\pi_{j}}{\pi_{j}-1}$ $=\prod_{j=1}^{k-1}\left(1+\frac{1}{\left\lceil cj\right\rceil}\right)\leq\exp\left(\sum_{j=1}^{k-1}\ln\left(1+\frac{1}{cj}\right)\right)$
$\leq\exp\left(\sum_{j=1}^{k-1}\frac{1}{cj}\right)\leq\exp\left(\frac{1}{c}+\int_{1}^{k}\frac{dx}{cx}\right)$
$=\exp\left(\frac{1}{c}+\frac{1}{c}\ln k\right)=(ek)^{1/c}.$ (73)
In consequence, we obtain these bounds
$f_{k}$ $=\prod_{j=1}^{k-1}\frac{\pi_{j}-1}{\pi_{j}}\leq\left(\frac{c+2}{ck+2}\right)^{1/c}\leq\left(\frac{c+2}{c}\right)^{1/c}k^{-1/c},$ (74)
$p_{k}$ $=f_{k}-f_{k+1}\leq\frac{1}{ck+1}\left(\frac{c+2}{ck+2}\right)^{1/c}\leq\frac{1}{c}\left(\frac{c+2}{c}\right)^{1/c}k^{-(c+1)/c},$ (75)
$w_{k}$ $=\sum_{j=2}^{k}\frac{1}{f_{j}}+\frac{\pi_{k}}{f_{k}}\leq\sum_{j=2}^{k}(ej)^{1/c}+(ck+2)(ek)^{1/c}$
[…]
$=\left(\frac{ck}{c+1}+ck+3\right)(ek)^{1/c}\leq\left(c+\frac{c}{c+1}+3\right)e^{1/c}k^{(c+1)/c}.$ (76)
Hence we may bound further
$u_{t}$ $=\min\left\{k:w_{k}>t\right\}\geq\left(\frac{t}{\left(c+\frac{c}{c+1}+3\right)e^{1/c}}\right)^{c/(c+1)},$ (77)
$g_{t}$ $=\max\left\{k:p_{k}^{-1}\leq t\right\}\leq\left(\frac{t}{c}\left(\frac{c+2}{c}\right)^{1/c}\right)^{c/(c+1)}.$ (78)
We also need the converse bounds
$p_{k}$ $\geq\frac{1}{ck+2}(ek)^{-1/c}\geq\frac{e^{-1/c}}{c+2}k^{-(c+1)/c},$ (79)
$g_{t}$ $\leq\left(\frac{te^{-1/c}}{c+2}\right)^{c/(c+1)}$ (80)
so as to derive
$q_{t}$ $=f_{g_{t}+1}\leq\left(\frac{c+2}{c}\right)^{1/c}\left(\frac{te^{-1/c}}{c+2}\right)^{-1/(c+1)}.$ (81)
As a result, bound $u_{t}-1\leq\operatorname{\mathbf{E}}V_{t}\leq g_{t}+tq_{t}$ implies the Herdan-Heaps law
$C_{1}t^{c/(c+1)}\leq\operatorname{\mathbf{E}}V_{t}\leq C_{2}t^{c/(c+1)}$ (82)
for certain constants $C_{1},C_{2}>0$. By Theorem 14, to derive Hilberg‘s law (7) with $\beta=c/(c+1)$, it suffices to show that $\operatorname{\mathbf{E}}K_{t}^{1/q}<\infty$ for some $q\geq 1$. In fact, for any $q>c$, we obtain
[…]
Figure 1: The number of types $V_{t}$ as a function of sequence length $t$ for periods $\pi_{k}=1+\left\lceil ck\right\rceil$ and seeds $\sigma_{k}=1$. The lines are theoretical predictions $V_{t}\sim t^{c/(c+1)}$, where the proportionality constant is chosen by least squares.
To provide some visualization of these results, in Figure 1, we present the growth of the actual number of types $V_{t}$ for three multiperiodic sequences with periods $\pi_{k}=1+\left\lceil ck\right\rceil$ and seeds $\sigma_{k}=1$, where $c\in\left\{0.5,1.0,2.0\right\}$. We notice a good agreement with the theoretical prediction $\operatorname{\mathbf{E}}V_{t}\sim t^{c/(c+1)}$.
6 Conclusion
Our toy example of multiperiodic processes extends mathematical analysis of Hilberg‘s law—the power law growth of redundancy for certain stationary sources, possibly entailing statistical language models [2, 3]. This topic is relevant not only for quantitative linguistics [4, 5] but also for large language models in machine learning, via the neural scaling law [6, 7, 8]. As it was shown in the previous research revolving around the example of Santa Fe processes [4, 5, 18, 10], the subject can be linked to traditional analyses of Zipf‘s law [11, 12]. This study has supplied further findings in this vein.
Prior versions
This work stems from rethinking and heavily editing an imperfect manuscript [29], which was posted to ArXiv on February 17, 2023. That manuscript has not been posted or published anywhere else but was presented as a seminar talk at the Math Machine Learning seminar in the Max Planck Institute for Mathematics in the Sciences.
Funding
This work received no external funding.
Acknowledgment
I thank Jan Mielniczuk, Szymon Jaroszewicz, Marcus Hutter, and Guido Montúfar for discussing prior versions of the paper.