### v4

# Multiperiodic Processes: Ergodic Sources with a Sublinear Entropy

Multiperiodic Processes:
Ergodic Sources with a Sublinear Entropy
Łukasz Dębowski Łukasz Dębowski is with the Institute of Computer Science, Polish Academy of Sciences, ul. Jana Kazimierza 5, 01-248 Warszawa, Poland, e-mail: ldebowsk@ipipan.waw.pl.
We construct multiperiodic processes—a simple example of stationary ergodic (but not mixing) processes over natural numbers that enjoy the vanishing entropy rate under a mild condition. Multiperiodic processes are supported on randomly shifted deterministic sequences called multiperiodic sequences, which can be efficiently generated using an algorithm called the Infinite Clock. Under a suitable parameterization, multiperiodic sequences exhibit relative frequencies of particular numbers given by Zipf‘s law. Exactly in the same setting, the respective multiperiodic processes satisfy an asymptotic power-law growth of block entropy, called Hilberg‘s law. Hilberg‘s law is deemed to hold for statistical language models, in particular.
Key words: ergodic processes, periodic sequences, entropy rate, Zipf‘s law, Hilberg‘s law
MSC 2020: 60G10, 62M20, 94A17
1 Introduction
Supplementing prior instances such as Santa Fe processes introduced in [1, 2], the aim of the present paper is to construct further examples of stationary sources over a countable alphabet that obey a power-law growth of block entropy, called Hilberg‘s law [3]. The processes constructed in this paper are called multiperiodic processes. Multiperiodic processes are stationary ergodic (but not mixing) and enjoy the vanishing entropy rate—in contrast to Santa Fe processes [1, 2], which are non-ergodic and usually have a positive entropy rate. As it has been demonstrated in [4], Hilberg‘s law—deemed to hold for natural language [3]—implies the better known neural scaling law, which is a power-law decay of the cross entropy rate and empirically holds for large language models trained on internet-sized corpora of texts [5, 6, 7, 8]. We hope that the toy examples of Santa Fe and multiperiodic processes may be useful as baseline artificial data sources to be compared with natural texts in quantitative linguistics and machine learning.
We cast our contribution against the background of previous research.
Preliminaries.
The notations to state our problem are as follows. Blocks of random variables are denoted $X_{j}^{k}:=(X_{j},X_{j+1},\ldots,X_{k})$. Let $H(X):=\operatorname{\mathbf{E}}\left(-\log P(X)\right)$ be the Shannon entropy of a discrete random variable $X$ where $\log$ is the base 2 logarithm, let H(X|Y):=H(X,Y)H(Y)H(X|Y):=H(X,Y)-H(Y) be the conditional entropy, and let I(X;Y):=H(X)H(X|Y)I(X;Y):=H(X)-H(X|Y) be the mutual information. Relation $\sim$ is defined as
$f(n)\sim g(n)\stackrel{{\scriptstyle\text{def}}}{{\iff}}\lim_{n\to\infty}\frac{\log f(n)}{\log g(n)}=1.$ (1)
For a stationary process $(X_{t})_{t\in\mathbb{Z}}$, we may define two limits
$h$ $:=\lim_{t\to\infty}\frac{H(X_{1}^{t})}{t}=\lim_{t\to\infty}H(X_{1}|X_{-t+1}^{0}),$ (2)
$E$ $:=\lim_{t\to\infty}\left(H(X_{1}^{t})-ht\right)=\lim_{t\to\infty}I(X_{-t+1}^{0};X_{1}^{t}).$ (3)
Limit $h$ is called entropy rate [9, 10], whereas $E$ is called excess entropy [11] or predictive information [12, 13, 14]. While entropy rate is a measure of process unpredictability, excess entropy is an intuitive measure of process memory. Although excess entropy only measures the memory capacity, without specifying how exactly the process future depends on the past, it can be given interesting interpretations. Some attempts were also made to generalize excess entropy to two-dimensional random fields [15, 16].
Memory conditions.
Conditions $E<\infty$ and $E=\infty$ have been studied in probability theory and its applications for quite a time. The oldest systematic results seem to concern Gaussian processes. In particular, Grenander and Szegő [17, Section 5.5] presented an integral formula for excess entropy (in disguise). Finch [18] evaluated this formula for autoregressive moving average (ARMA) processes, which turned out to yield $E<\infty$. More generally, we have $E<\infty$ if and only if $\sum_{k=1}^{\infty}k\alpha_{k}^{2}<\infty$ and $\left|\alpha_{k}\right|<1$ for the partial autocorrelation $\alpha_{k}$ [19]. In particular, under a positive and continuous spectral density, Li [20] showed that $E<\infty$ if and only if the autocorrelation $\rho_{k}:=\operatorname{Corr}(X_{0};X_{k})$ satisfies $\sum_{k=1}^{\infty}k\rho_{k}^{2}<\infty$.
Let us proceed to the countable alphabet. It can be easily shown, by the data-processing inequality, that condition $E<\infty$ holds for hidden Markov processes with a finite number of states [11]. Mahoney, Ellison and Crutchfield [21, 22] developed a formula for excess entropy in terms of predictive and retrodictive $\epsilon$-machines, which are minimal unifilar hidden Markov representations of the process [23, 24]. Travers and Crutchfield [25] and Dębowski [26] exhibited examples of hidden Markov processes with a countably infinite number of hidden states that obey $E=\infty$ and the further discussed Hilberg‘s law (4). A few other older papers concerned also isolated examples of processes with $E=\infty$ over a finite alphabet. For instance, Bradley [27] constructed the first example of a mixing process having this property. Gramss [28] investigated a process which is formed by the frequencies of 0‘s and 1‘s in the rabbit sequence.
A more systematic study of condition $E=\infty$ for the countable alphabet case was inspired by the publication by Hilberg [3] who redrew the plot by Shannon [29] in the log-log scale and proposed that relationship
$H(X_{1}^{t})-ht\sim t^{\beta},$ (4)
where $\beta\in(0,1)$, holds for natural language. Although the first reception of Hilberg‘s law (4) came from physicists [30, 31, 13, 11, 24], there is a line of research that seeks to combine a rigorous information-theoretic approach to Hilberg‘s law with general linguistic insights [32, 1, 2, 33, 34, 35] such as Zipf‘s law in particular [36, 37]. The basic observation is that condition $E=\infty$ holds if the process is strongly non-ergodic in the sense that its shift-invariant algebra $\mathcal{I}$ is non-atomic. It is so because
$E=H(\mathcal{I})+I(X_{-\infty}^{0};X_{1}^{\infty}|\mathcal{I})$ (5)
and $H(\mathcal{A})=\infty$ for a non-atomic algebra $\mathcal{A}$ under an appropriate generalization of Shannon information measures to arbitrary algebras of events [38, 39, 40, 41, 1, 42].
Santa Fe processes.
The ergodic decomposition of excess entropy (5) inspires a family of linguistically motivated examples of processes that obey Hilberg‘s law (4). These examples are called Santa Fe processes [1, 2] and were independently considered by Hutter [43]. We recall this family of examples at length for their certain structural similarities and differences to the multiperiodic processes constructed in this paper. The Santa Fe processes come in a number of varieties—some are strongly non-ergodic, some are ergodic. Their common simple idea is to decompose each token $X_{t}$ as a pair of a natural number $K_{t}$ and an additional bit—which is copied from a certain sequence of bits $(Z_{k})_{k\in\mathbb{N}}$ by taking the bit at position $K_{t}$. In other words, each text token $X_{t}$ may be written as a pair
$X_{t}=(K_{t},Z_{K_{t}}),$ (6)
where $(Z_{k})_{k\in\mathbb{N}}$ is the sequence of bits and $(K_{t})_{t\in\mathbb{Z}}$ is the sequence of numbers.
The Santa Fe decomposition (6) is interesting for its semantic interpretation, discussed in [2, 34]. We may interpret that pairs $(K_{t},Z_{K_{t}})$ are logical statements that describe a certain binary sequence $(Z_{k})_{k\in\mathbb{N}}$ at random but in a non-contradictory way. Namely, if statements $(k,Z_{k})$ and $(k^{\prime},Z_{k^{\prime}})$ describe the same bit $(k=k^{\prime})$ then they assert the same value $(Z_{k}=Z_{k^{\prime}})$. Thus, we may interpret that sequence $(Z_{k})_{k\in\mathbb{N}}$ expresses some unbounded immutable knowledge which is only partially accessed and described in finite texts. For this reason, bits $Z_{k}$ appearing in representation (6) were called the ’’(elementary) facts‘‘ [2], whereas sequence $(K_{t})_{t\in\mathbb{Z}}$ was called the ’’narration (process)‘‘ [34]. Thus, we may say that formula (6) decomposes a corpus of texts $(X_{t})_{t\in\mathbb{Z}}$ into ’’knowledge‘‘ $(Z_{k})_{k\in\mathbb{N}}$ and ’’narration‘‘ $(K_{t})_{t\in\mathbb{Z}}$. To complete the naming, individual variables $K_{t}$ might be called ’’topics‘‘.
There may exist connections between the Santa Fe decomposition (6) and algorithmic information theory [44], see also [34, Chapter 1]. In the role of immutable knowledge, the random probability $Y:=\sum_{k=1}^{\infty}2^{-k}Z_{k}\in(0,1)$ resembles Chaitin‘s number $\Omega\in(0,1)$, also called the halting probability [45, 46]. The number $\Omega$ is algorithmically random [47] and encodes a large body of mathematical knowledge in a dense form. However, equating the values of probabilities $Y$ and $\Omega$ is preposterous for two reasons. The first of them, a formal one, is that $Y$ is a random variable, whereas $\Omega$ is a fixed number. The second reason, philosophically more important, is that $\Omega$ does not seem computable in a sufficiently strong sense to arise as an effect of natural processes [48]. If there are some links between the Santa Fe decomposition (6) and algorithmic information theory, they must be more subtle.
The default example of Santa Fe processes assumes that narration $(K_{t})_{t\in\mathbb{Z}}$ is a memoryless source with the marginal Zipf distribution
$P(K_{t}=k)\sim k^{-\alpha},$ (7)
where $\alpha>1$ [36, 37]. Moreover, it was assumed that knowledge $(Z_{k})_{k\in\mathbb{N}}$ is a sequence of independent fair coin flips, independent of process $(K_{t})_{t\in\mathbb{Z}}$. Under these conditions, we obtain Hilberg‘s law (4) with $\beta=1/\alpha$. The reason for this law is as follows. Denote the random set of types [^1] We adopt the distinction of terms “type” and “token” borrowed from quantitative linguistics [49, 50]: A token is an element of a sequence with possible repetitions, whereas a type is an element of a set, which excludes repetitions.
$\mathcal{V}_{t}:=\left\{K_{1},K_{2},\ldots,K_{t}\right\}.$ (8)
For memoryless sources with Zipf‘s distribution (7), under mild conditions [51, 52], we have Heaps‘ law
$\operatorname{\mathbf{E}}\operatorname{\#}\mathcal{V}_{t}\sim t^{1/\alpha}$ (9)
[49, 53, 50]. Consequently, we may express
$H(X_{1}^{t})-H(K_{1}^{t})=H(\left\{(k,Z_{k}):k\in\mathcal{V}_{t}\right\}|K_{1}^{t})=\operatorname{\mathbf{E}}\operatorname{\#}\mathcal{V}_{t}\sim t^{1/\alpha},$ (10)
which implies $H(K_{1}^{t})=H(K_{0})t=ht$, thus recovering Hilberg‘s law (4).
The above described Santa Fe process is non-ergodic and takes values in a countably infinite alphabet. The reason for its non-ergodicity is that coin flips $(Z_{k})_{k\in\mathbb{N}}$ are measurable with respect to the shift-invariant algebra of process $(X_{t})_{t\in\mathbb{Z}}$. Thus each particular value of sequence $(Z_{k})_{k\in\mathbb{N}}$ defines a distinct ergodic component [1, 34]. However, replacing individual coin flips $Z_{k}$ with slowly evolving Markov chains $(Z_{kt})_{t\in\mathbb{Z}}$, we obtain processes which are (strongly) mixing hence ergodic [33]. Coding natural numbers as binary strings, we obtain processes over a finite alphabet [2, 35]. All these examples enjoy Hilberg‘s law (4) with a positive rate $h>0$.
Further examples.
To complete the collection of invented sources, we have been interested in finding an example of a process with Hilberg‘s law (4) and the vanishing Shannon entropy rate, $h=0$. In [54, 34], we have constructed the random hierarchical association processes (RHA), which satisfy (4) with $h=0$ but are probably non-ergodic. The construction of RHA processes is so complicated that we do not reproduce it here. In the end, we desired a simpler construction that would be also ergodic.
Our curiosity in such asymptotically deterministic models was initially driven by the observation that the maximal repetition length for natural language obeys a cube logarithmic growth [55]. Consequently, the conditional Rényi entropy rate may vanish for natural language [56, 57]. Later, our interest was reinforced by the research in the neural scaling law [5, 6, 7, 8], which suggested that Hilberg‘s law (4) holds for large language models with $h=0$. We are agnostic about the actual value of particular entropy rates for natural language but we have wished to better understand the mathematical construction of zero entropy rate processes.
Present contribution.
This paper supplies another toy example called multiperiodic processes. Multiperiodic processes $(K_{t})_{t\in\mathbb{Z}}$ are asymptotically deterministic sequences of random natural numbers. They are not mixing but they are ergodic and, under a certain parameterization, they satisfy Hilberg‘s law with the vanishing Shannon entropy rate, namely,
$H(K_{1}^{t})\sim t^{\beta}.$ (11)
The essential idea of these processes is to define a probability measure on a set of efficiently computable but randomly shifted deterministic sequences, called multiperiodic sequences. Under a certain parameterization, the relative frequencies of particular numbers in these sequences approach Zipf‘s distribution (7). This suffices to imply Hilberg‘s law (11) with exponent $\beta=1/\alpha$ by a reasoning similar to bound (10). Namely, for the set of types (8), we also have Heaps‘ law (9) and we may express
$H(K_{1}^{t})=H(\left\{(k,\Sigma_{k}):k\in\mathcal{V}_{t}\right\})\sim\operatorname{\mathbf{E}}\operatorname{\#}\mathcal{V}_{t}\sim t^{1/\alpha},$ (12)
where $(\Sigma_{k})_{k\in\mathbb{N}}$ are certain random variables called seeds. The seeds parameterize random shifts of the underlying deterministic multiperiodic sequences. Variables $(\Sigma_{k})_{k\in\mathbb{N}}$ are probabilistically independent and measurable with respect to process $(K_{t})_{t\in\mathbb{Z}}$ but are not measurable with respect to the shift-invariant algebra of process $(K_{t})_{t\in\mathbb{Z}}$. Thus the multiperiodic process $(K_{t})_{t\in\mathbb{Z}}$ can be ergodic. Moreover, we do not need to extend $(K_{t})_{t\in\mathbb{Z}}$ with additional coin flips $(Z_{k})_{k\in\mathbb{N}}$ and the Santa Fe construction (6) to obtain Hilberg‘s law. This remark concludes our brief presentation of the main results.
Organization of the article.
The organization of the paper is as follows. In Section 2, we define two objects: multiperiodic sequences in Section 2.1 and multiperiodic processes in Section 2.2. We analyze their statistics in Section 3: the relative frequencies in Section 3.1, the waiting times in Section 3.2, and the number of types in Section 3.3. In Section 4, we inspect the flow of information: estimation of seeds given a finite sample in Section 4.1 and the block entropy in Section 4.2. These general developments are supplemented by two examples in Section 5: constant periods in Section 5.1 and linear periods in Section 5.2. Section 6 concludes the article.
2 Definitions
In this section, we construct two kinds of objects. The first one are multiperiodic sequences and the second one are muliperiodic processes. We also prove first theorems concerning these constructions: time reversal symmetry, periodicity of decimation, existence conditions, stationarity, and ergodicity. These properties will be used later for bounding some quantities of interest.
2.1 Multiperiodic sequences
Multiperiodic sequences are an idea which we have originally generalized from work [58]. The authors only considered a fixed multiperiodic sequence with particular periods and seeds—to state a certain negative result in universal prediction. Isolated instances of multiperiodic sequences are noted in the on-line encyclopedia of integer sequences [59] and can be connected to other problems, such as Mancala-type games [60]. Distilling the general construction, especially considering the idea of random seeds that matters in the probabilistic context, seems our contribution.
Definition 1 (multiperiodic sequence).
Consider a sequence $(\pi_{k})_{k\in\mathbb{N}}$ of natural numbers $\pi_{k}\geq 2$. Let $(\sigma_{k})_{k\in\mathbb{N}}$ be another sequence of natural numbers $\sigma_{k}\in\left\{1,2,\ldots,\pi_{k}\right\}$ such that set $\left\{k\in\mathbb{N}:\sigma_{k}=1\right\}$ is infinite. Parameters $\pi_{k}$ are called periods, whereas parameters $\sigma_{k}$ are called seeds. Given $(\pi_{k})_{k\in\mathbb{N}}$ and $(\sigma_{k})_{k\in\mathbb{N}}$, the multiperiodic sequence $(k_{t})_{t\in\mathbb{Z}}$ is the result of the initial condition
$\phi_{k,1}=\sigma_{k}$ (13)
and recursion
$k_{t}$ $=\min\left\{k\in\mathbb{N}:\phi_{k,t}=1\right\},$ (14)
$\phi_{k,t+1}$ $=\begin{cases}\phi_{k,t}-1,&k<k_{t},\\ \pi_{k},&k=k_{t},\\ \phi_{k,t},&k>k_{t},\end{cases}$ (15)
which can be inverted as
$k_{t}$ $=\min\left\{k\in\mathbb{N}:\phi_{k,t+1}=\pi_{k}\right\},$ (16)
$\phi_{k,t}$ $=\begin{cases}\phi_{k,t+1}+1,&k<k_{t},\\ 1,&k=k_{t},\\ \phi_{k,t+1},&k>k_{t}.\end{cases}$ (17)
The auxiliary matrix $(\phi_{k,t})_{k\in\mathbb{N},t\in\mathbb{Z}}$ is called the hand matrix.
Here we present two examples of multiperiodic sequences.
Example 1 (A001511 in [59]).
Consider the multiperiodic sequence $(k_{t})_{t\in\mathbb{N}}$ with constant periods $\pi_{k}=2$ and seeds $\sigma_{k}=1$, namely,
$1,2,1,3,1,2,1,4,1,2,1,3,1,2,1,5,1,2,1,3,1,2,1,4,1,2,1,\ldots$ (18)
Equivalently, $k_{t}$ is the number of 22‘s dividing $2t$. The equivalence stems from the observation that $k_{t}$ is the position in the binary expansion of $t$ that is incremented when increasing from t1t-1 to $t$. This sequence was applied by Kalociński and Steifer [58] for a negative result in universal prediction.
Our applications are closer to this example:
Example 2 (A028920 in [59]).
Let $\pi_{k}=1+k$ and $\sigma_{k}=1$ for all $k\in\mathbb{N}$. Then the multiperiodic sequence $(k_{t})_{t\in\mathbb{N}}$ is
$1,2,1,3,1,4,1,2,1,5,1,6,1,2,1,3,1,7,1,2,1,8,1,4,1,2,1,\ldots$ (19)
This sequence was discussed by Broline and Loeb [60] in the context of Mancala-type games. The algorithm for generating this sequence is isomorphic with recursion (13)–(17).
As we show in Section 5.2, for periods $\pi_{k}\approx ck$, the relative frequency of number $k$ approaches Zipf‘s law (7) with $\alpha=(c+1)/c$.
Recursion (13)–(17) for $k\in\mathbb{N}$ can be rewritten as Algorithm 1, called the Infinite Clock. The Infinite Clock operates with infinite data structures for simplicity of presentation. It works like a clock with infinitely many hands—indexed by natural numbers—that move counterclockwise, hence the name. As in the usual clock, each hand moves with a different speed, dictated by its specific period. The initial positions of hands are given by seeds. The algorithm outputs the value of the hand index whenever a hand passes noon.
Algorithm 1 The Infinite Clock
1:Sequence of $\pi_{k}\in\mathbb{N}$ for $k\in\mathbb{N}$. $\triangleright$ The periods.
2:Sequence of $\phi_{k}:=\sigma_{k}\in\left\{1,2,\ldots,\pi_{k}\right\}$ for $k\in\mathbb{N}$. $\triangleright$ The hands.
3:Sequence of $k_{t}\in\mathbb{N}$ for $t\in\mathbb{N}$. $\triangleright$ The multiperiodic sequence.
4:for $t\in\mathbb{N}$ do
5:  $k^{*}:=0$
6:  $k:=1$
7:  while $k^{*}=0$ do
8:   if $\phi_{k}>1$ then
9:     $\phi_{k}:=\phi_{k}-1$
10:   else
11:     $k^{*}:=k$
12:   $k:=k+1$
13:  $\phi_{k^{*}}:=\pi_{k^{*}}$
14:  $k_{t}:=k^{*}$
To be precise, Algorithm 1 operates with hand variables $(\phi_{k})_{k\in\mathbb{N}}$, where each hand $\phi_{k}$ takes values in set $\left\{1,2,\ldots,\pi_{k}\right\}$ and is initialized with the value of the corresponding seed, $\phi_{k}:=\sigma_{k}$. To determine the value of the next element $k_{t}$ of the sequence, the algorithm decrements consecutive hands $(\phi_{k})_{k\in\mathbb{N}}$ starting from index $k=1$ until it finds an index $k^{*}$ such that $\phi_{k^{*}}=1$. Then the loop is broken, this particular hand is reset to its own maximal value $\phi_{k^{*}}:=\pi_{k^{*}}$, and the corresponding index $k^{*}$ of the hand is output as the next element of the sequence $k_{t}:=k^{*}$.
Let us observe that the computation falls into an infinite loop in lines 4–9 of Algorithm 1 if and only if there is no $k\in\mathbb{N}$ such that $\phi_{k}=1$. This happens precisely if and only if set $\left\{k\in\mathbb{N}:\sigma_{k}=1\right\}$ is finite, which is excluded by Definition 1. The easiest way to prevent this situation is to set the minimal seeds $\sigma_{k}=1$ for all $k\in\mathbb{N}$.
As a general fact, we notice a time-reversal symmetry.
Theorem 1.
For the multiperiodic sequence $(k_{t})_{t\in\mathbb{Z}}$ with periods $\pi_{k}$ and seeds $\sigma_{k}$, sequence $(k_{1-t})_{t\in\mathbb{Z}}$ is the multiperiodic sequence with periods $(\pi_{k})_{k\in\mathbb{N}}$ and seeds $(\sigma^{R}_{k})_{k\in\mathbb{N}}$ defined as $\sigma^{R}_{k}:=\pi_{k}-\sigma_{k}+1$.
Proof.
Let $(\phi^{R}_{k,t})_{k\in\mathbb{N},t\in\mathbb{Z}}$ and $(k^{R}_{t})_{t\in\mathbb{Z}}$ be defined analogously as (14)–(17) but with the initial condition $\phi^{R}_{k,1}=\sigma^{R}_{k}$. Then by induction we obtain $\phi^{R}_{k,t}=\pi_{k}-\phi_{k,1-t}+1$ and $k^{R}_{t}=k_{1-t}$. ∎
Moreover, we observe a partial periodicity of decimated multiperiodic sequences, which are also multiperiodic.
Theorem 2.
Let $(k^{\geq r}_{t})_{t\in\mathbb{N}}$ be the subsequence of the multiperiodic sequence $(k_{t})_{t\in\mathbb{N}}$ with periods $\pi_{k}$ and seeds $\sigma_{k}$ from which we have removed all tokens $k_{t}<r$. We claim that:
  1. 1.
    Sequence $(k^{\geq r}_{t}-r+1)_{t\in\mathbb{N}}$ is the multiperiodic sequence with periods $\pi^{\geq r}_{k}=\pi_{k-r+1}$ and seeds $\sigma^{\geq r}_{k}=\sigma_{k-r+1}$.
  2. 2.
    We have $k^{\geq r}_{t}=r\iff t\equiv\sigma_{r}\mod\pi_{r}$.
Proof.
We proceed by induction on $r$. As for the initial step, we have $(k^{\geq 1}_{t})_{t\in\mathbb{N}}=(k_{t})_{t\in\mathbb{N}}$. Thus claim 1. holds by definition, whereas claim 2. follows since $k_{t}=1$ if and only if $t\equiv\sigma_{1}\mod\pi_{1}$. As for the inductive step, it is sufficient to show that claim 1. holds, whereas claim 2. follows by claim 1. for the inductive step combined with claim 2. for the initial step. The proof of claim 1. for the inductive step rests on the fact that to obtain $(k^{\geq r}_{t})_{t\in\mathbb{N}}$ instead of $(k_{t})_{t\in\mathbb{N}}$, it suffices to initialize $k:=r$ instead of $k:=1$ in line 3 of Algorithm 1. ∎
2.2 Multiperiodic processes
Multiperiodic sequences can be applied to define stationary ergodic processes. The basic idea is to apply a random initialization, that is, uniformly random seeds. This leads to the next construction, called a multiperiodic process. We observe that the existence of the process is conditional—given that the value $1$ appears in the seed sequence infinitely often almost surely.
Definition 2 (multiperiodic process).
The multiperiodic process $(K_{t})_{t\in\mathbb{Z}}$ with periods $(\pi_{k})_{k\in\mathbb{N}}$ is the random multiperiodic sequence with periods $(\pi_{k})_{k\in\mathbb{N}}$ and probabilistically independent random seeds $(\Sigma_{k})_{k\in\mathbb{N}}$ with uniform distributions
$P(\Sigma_{k}=\sigma_{k})=\frac{1}{\pi_{k}},\quad\sigma_{k}\in\left\{1,2,\ldots,\pi_{k}\right\}$ (20)
provided set $\left\{k\in\mathbb{N}:\Sigma_{k}=1\right\}$ is infinite almost surely.
Applying the complementary Borel-Cantelli lemmas for independent events, we derive a simple criterion of existence of a multiperiodic process.
Theorem 3.
The multiperiodic process exists if and only if its periods satisfy
$\sum_{k=1}^{\infty}\frac{1}{\pi_{k}}=\infty.$ (21)
Proof.
Observe that events $(\Sigma_{k}=1)$ are independent. Hence by the Borel-Cantelli lemmas [61, Theorems 4.3 and 4.4], set $\left\{k\in\mathbb{N}:\Sigma_{k}=1\right\}$ is infinite almost surely if and only if
$\sum_{k=1}^{\infty}P(\Sigma_{k}=1)=\infty.$ (22)
In view of the uniform distributions of $(\Sigma_{k})_{k\in\mathbb{N}}$, we obtain the claim. ∎
Moreover, through the approximation by periodic processes, each multiperiodic process admits a frequency interpretation of its probability distribution. Namely, it inherits the properties of being stationary and ergodic from the approximating periodic sources.
Theorem 4.
The multiperiodic process $(K_{t})_{t\in\mathbb{Z}}$ is stationary and ergodic but is not mixing.
Proof.
A sequence $(a_{t})_{t\in\mathbb{Z}}$ is called $p$-periodic if
$(a_{t+s})_{t\in\mathbb{Z}}=(a_{t})_{t\in\mathbb{Z}}\iff s\equiv 0\mod p.$ (23)
A stochastic process $(A_{t})_{t\in\mathbb{Z}}$ is called $p$-periodic if
$P((A_{t})_{t\in\mathbb{Z}}=(a_{t+s})_{t\in\mathbb{Z}})=\frac{1}{p},\quad s\geq 0,$ (24)
for a certain $p$-periodic sequence $(a_{t})_{t\in\mathbb{Z}}$. It is known that $p$-periodic processes are stationary and ergodic but they are not mixing [34, Section 4.1].
Let an $m\in\mathbb{N}$. Let $K^{\leq m}_{t}:=\min\left\{K_{t},m\right\}$. Observe that process $(K^{\leq m}_{t})_{t\in\mathbb{Z}}$ depends only on seeds $(\Sigma_{k})_{k\leq m}$. Denote $p_{m}:=\prod_{k=1}^{m}\pi_{k}$. We will show that process $(K^{\leq m}_{t})_{t\in\mathbb{Z}}$ is $p$-periodic where $p\leq p_{m}$.
Let $(\Phi_{k,t})_{k\in\mathbb{N},t\in\mathbb{Z}}$ be the hand matrix of sequence $(K_{t})_{t\in\mathbb{Z}}$. By recursion (14)–(17), there is a bijection $(\Phi_{k,t})_{k\leq m}\leftrightarrow(\Phi_{k,t+1})_{k\leq m}$. Hence process $(\Phi_{k,t})_{k\leq m,t\in\mathbb{Z}}$ is a Markov process over time $t$. There are exactly $p_{m}$ distinct states of hands $(\phi_{k})_{k\leq m}$ and it can be shown that all states communicate. Hence Markov process $(\Phi_{k,t})_{k\leq m,t\in\mathbb{Z}}$ is irreducible, each state $(\phi_{k})_{k\leq m}$ is visited exactly once within the period $p_{m}$, and the uniform marginal distribution of the initial state $(\Phi_{k,1})_{k\leq m}=(\Sigma_{k})_{k\leq m}$ is the invariant distribution. In turn, process $(\Phi_{k,t})_{k\leq m,t\in\mathbb{Z}}$ is stationary, ergodic, and $p_{m}$-periodic. Since each $K^{\leq m}_{t}$ is a function of $(\Phi_{k,t})_{k\leq m}$, hence process $(K^{\leq m}_{t})_{t\in\mathbb{Z}}$ is $p$-periodic where $p\leq p_{m}$.
Thus process $(K^{\leq m}_{t})_{t\in\mathbb{Z}}$ is stationary and ergodic but is not mixing. Since $m\in\mathbb{N}$ is arbitrarily chosen, hence also process $(K_{t})_{t\in\mathbb{Z}}$ is stationary and ergodic but is not mixing. The reason is that verifying stationarity, ergodicity, and mixing can be reduced to checking probabilities of finite blocks by [34, Theorem 4.10] or [62, Lemma 7.15]. ∎
3 Statistics
In this section, we analyze several statistics of multiperiodic sequences and multiperiodic processes. We begin with the relative frequencies and the marginal distribution. The waiting times are the second kind of statistics. The third sort of investigated quantities relate to the number of observed types. Some results concerning these quantities are established by applying the decimation properties shown in Theorem 2, the Birkhoff ergodic theorem, the union bound, and the Borel-Cantelli lemma.
3.1 Relative frequencies
By the decimation property established in Theorem 2, for a multiperiodic sequence there exist limiting relative frequences.
Theorem 5.
Let $(k_{t})_{t\in\mathbb{Z}}$ be a multiperiodic sequence with periods $(\pi_{k})_{k\in\mathbb{N}}$. The relative frequencies exist and equal
$\lim_{t\to\infty}\frac{1}{t}\sum_{j=1}^{t}{\mathbf{1}}{\left\{k_{j}\geq k\right\}}=f_{k}:=\prod_{j=1}^{k-1}\frac{\pi_{j}-1}{\pi_{j}}.$ (25)
Proof.
Let $(k_{t}^{\geq r})_{t\in\mathbb{N}}$ be the subsequence of $(k_{t})_{t\in\mathbb{N}}$ from which we have removed all tokens $k_{t}<r$. We proceed by induction on $k$. First, we observe
$\lim_{t\to\infty}\frac{1}{t}\sum_{j=1}^{t}{\mathbf{1}}{\left\{k_{j}\geq 1\right\}}=1=f_{1}.$ (26)
Second, in view of claim 2. of Theorem 2, we notice
$\lim_{t\to\infty}\frac{1}{t}\sum_{j=1}^{t}{\mathbf{1}}{\left\{k_{j}\geq k\right\}}$
$=\left(\lim_{t\to\infty}\frac{1}{t}\sum_{j=1}^{t}{\mathbf{1}}{\left\{k_{j}\geq k-1\right\}}\right)\left(\lim_{t\to\infty}\frac{1}{t}\sum_{j=1}^{t}{\mathbf{1}}{\left\{k_{j}^{(k-1)}\geq k\right\}}\right)$
$=f_{k-1}\left(\frac{\pi_{k-1}-1}{\pi_{k-1}}\right)=f_{k}.$ (27)
Hence the claim follows. ∎
Hence by ergodicity of the multiperiodic process, the limiting relative frequencies are equal to the marginal distribution of the process.
Theorem 6.
The multiperiodic process $(K_{t})_{t\in\mathbb{Z}}$ with periods $(\pi_{k})_{k\in\mathbb{N}}$ has the marginal distribution
$P(K_{0}=k)=p_{k}:=f_{k}-f_{k+1},$ (28)
where $f_{k}$ is the marginal distribution function (25).
Proof.
We obtain distribution (28) by the Birkhoff ergodic theorem, which establishes the almost sure equality of probabilities and relative frequencies for stationary ergodic processes [63, 64]. ∎
3.2 Waiting times
Another important statistic of a stochastic process is the waiting time. Following [65, 66], we define the waiting time as the first position in the process where the value of interest occurs. That is we put
$W_{k}:=\inf\left\{t\in\mathbb{N}:K_{t}=k\right\}.$ (29)
The celebrated Kac lemma [65] predicts that the conditional expectation of the waiting time equals the inverse probability of the recurring value. Namely, we have
$\operatorname{\mathbf{E}}\left(W_{k}|K_{0}=k\right)=\frac{1}{P(K_{0}=k)}$ (30)
for a stationary ergodic process $(K_{t})_{t\in\mathbb{Z}}$.
Complementing this result, for a multiperiodic process, we can obtain a uniform bound as follows.
Theorem 7.
For a multiperiodic process $(K_{t})_{t\in\mathbb{Z}}$ with periods $(\pi_{k})_{k\in\mathbb{N}}$, we have a uniform bound
$W_{k}\leq w_{k}:=\sum_{j=2}^{k}\frac{1}{f_{j}}+\frac{\pi_{k}}{f_{k}},$ (31)
where $f_{k}$ is the marginal distribution function (25).
Proof.
Consider a prefix $(K_{1},K_{2},\ldots,K_{T_{1}})$ of the multiperiodic process. When we delete all $K_{t}<r$, the sequence shortens to $(K^{\geq r}_{1},K^{\geq r}_{2},\ldots,K^{\geq r}_{T_{r}})$. We observe that $W_{k}\leq T_{1}$ if $T_{k}=\pi_{k}$, whereas the partial periodicity implies
$T_{r}-T_{r+1}\leq\frac{T_{r}}{\pi_{r}}+1.$ (32)
Hence we have
$T_{r}\leq(T_{r+1}+1)\cdot\frac{\pi_{r}}{\pi_{r}-1}$ (33)
and consequently we obtain
$W_{k}$ $\leq\left(\ldots\left(\left(\left(\pi_{k}+1\right)\frac{\pi_{k-1}}{\pi_{k-1}-1}+1\right)\frac{\pi_{k-2}}{\pi_{k-2}-1}+1\right)\ldots\right)\frac{\pi_{1}}{\pi_{1}-1}$
$=\sum_{j=2}^{k}\frac{1}{f_{j}}+\frac{\pi_{k}}{f_{k}}=w_{k}$ (34)
in view of identity
$\left(\ldots\left(\left(\left(a_{k}+1\right)a_{k-1}+1\right)a_{k-2}+1\right)\ldots\right)a_{1}=\sum_{j=1}^{k}\prod_{i=1}^{j}a_{i}.$ (35)
Moreover, for increasing periods, we obtain that the waiting times increase monotonically, whereas the marginal probabilities are monotonically decreasing.
Definition 3 (increasing periods).
Periods $(\pi_{k})_{k\in\mathbb{N}}$ are called increasing if $\pi_{k+1}\geq\pi_{k}$ for all $k\in\mathbb{N}$.
Theorem 8.
For increasing periods $(\pi_{k})_{k\in\mathbb{N}}$, we have:
$w_{k+1}$ $\geq w_{k},$ (36)
$P(K_{0}=k+1)$ $\leq P(K_{0}=k).$ (37)
Proof.
We have $w_{k+1}\geq w_{k}$ since
$\pi_{k}+1\leq\pi_{k+1}+1\leq(\pi_{k+1}+1)\frac{\pi_{k}}{\pi_{k}-1}+1.$ (38)
Similarly $P(K_{0}=k)=f_{k}-f_{k+1}\geq f_{k+1}-f_{k+2}=P(K_{0}=k+1)$ since
$1-\frac{\pi_{k}-1}{\pi_{k}}\geq\frac{\pi_{k}-1}{\pi_{k}}\left(1-\frac{\pi_{k+1}-1}{\pi_{k+1}}\right).$ (39)
3.3 Number of types
Now we proceed to bounding the number of types observed in a finite sample. We prove results of a varying generality. Some of these propositions hold for general stationary sources and may be useful for quantitative linguistic applications revolving around studying the type-token ratio [50].
Let us denote the random set of observed types
$\mathcal{V}_{t}:=\left\{K_{1},K_{2},\ldots,K_{t}\right\}=\left\{k\in\mathbb{N}:W_{k}\leq t\right\}.$ (40)
We denote its cardinality, the random number of types in a finite sample,
$V_{t}:=\operatorname{\#}\mathcal{V}_{t}.$ (41)
To bound the number of types effectively, we introduce four other statistics:
$u_{t}$ $:=\min\left\{k\in\mathbb{N}:w_{k}>t\right\},$ (42)
$U_{t}$ $:=\min\left\{k\in\mathbb{N}:k\not\in\mathcal{V}_{t}\right\},$ (43)
$g_{t}$ $:=\max\left\{k\in\mathbb{N}:P(K_{0}=k)\geq\frac{1}{t}\right\},$ (44)
$q_{t}$ $:=P(K_{0}>g_{t}).$ (45)
The first result is a simple sandwich bound for multiperiodic processes.
Theorem 9.
For a multiperiodic process $(K_{t})_{t\in\mathbb{Z}}$, we have
$u_{t}-1\leq U_{t}-1\leq V_{t}.$ (46)
Proof.
The inequalities follow by $w_{k}\geq W_{k}$ and $U_{t}-1\leq\operatorname{\#}\mathcal{V}_{t}$. ∎
The subsequent two statements pertain to general stationary processes. We note that functions $g_{t}$ and $q_{t}$ were introduced by Khmaladze [52] for memoryless sources. Generalizing his technique from memoryless processes to stationary ones, we can prove this proposition.
Theorem 10.
For a stationary process $(K_{t})_{t\in\mathbb{Z}}$ over natural numbers, we have
$\operatorname{\mathbf{E}}V_{t}\leq g_{t}+tq_{t}.$ (47)
Proof.
We observe
$\operatorname{\mathbf{E}}V_{t}=\operatorname{\mathbf{E}}\sum_{k=1}^{\infty}{\mathbf{1}}{\left\{k\in\mathcal{V}_{t}\right\}}=\sum_{k=1}^{\infty}P(k\in\mathcal{V}_{t})\leq g_{t}+\sum_{k>g_{t}}P(k\in\mathcal{V}_{t})$ (48)
and the union bound for stationary processes
$P(k\in\mathcal{V}_{t})=P(K_{1}=k\lor\ldots\lor K_{t}=k)\leq tP(K_{0}=k).$ (49)
Hence the claim follows. ∎
Quotient $V_{t}/t$ is called the type-token ratio in quantitative linguistics [49, 50]. It is easy to see that it tends to zero in the stationary case.
Theorem 11.
For a stationary process $(K_{t})_{t\in\mathbb{Z}}$ over natural numbers,
$\lim_{t\to\infty}\frac{\operatorname{\mathbf{E}}V_{t}}{t}$ $=0.$ (50)
Proof.
Applying the ideas from the proof of the previous theorem, we observe that bound
$\operatorname{\mathbf{E}}V_{t}\leq h_{t}+tP(K_{0}>h_{t})$ (51)
holds for any function $h_{t}$. In particular, for an $\epsilon>0$, we may take $h_{t}=\epsilon t/2$. For all sufficiently large $t$, we observe $P(K_{0}>h_{t})\leq\epsilon/2$. Hence, for these $t$, we have $\operatorname{\mathbf{E}}V_{t}/t\leq h_{t}/t+P(K_{0}>h_{t})\leq\epsilon$. By arbitrariness of $\epsilon$, we derive (50). ∎
We will need a stronger bound that holds under a mild condition.
Theorem 12.
For a stationary process $(K_{t})_{t\in\mathbb{Z}}$ over natural numbers such that $\operatorname{\mathbf{E}}(\log K_{0})^{3}<\infty$, we have
$\lim_{t\to\infty}\frac{\operatorname{\mathbf{E}}V_{t}\log t}{t}$ $=0,$ (52)
$\lim_{t\to\infty}\frac{V_{t}\log t}{t}$ $=0\text{ a.s.}$ (53)
Proof.
Observe by the Markov inequality that
$\sum_{k=1}^{\infty}kP\left(K_{0}\geq\frac{2^{k}}{k^{3}}\right)$ $\leq\sum_{k=1}^{\infty}kP\left((\log K_{0})^{3}\geq(k-3\log k)^{3}\right)$
$\leq\operatorname{\mathbf{E}}(\log K_{0})^{3}\sum_{k=1}^{\infty}\frac{k}{(k-3\log k)^{3}}<\infty.$ (54)
Choose $h_{t}=t/(\log t)^{3}$ in (51). Then we obtain
$\frac{\operatorname{\mathbf{E}}V_{2^{k}}k}{2^{k}}\leq\frac{1}{k^{2}}+kP\left(K_{0}\geq\frac{2^{k}}{k^{3}}\right).$ (55)
Consequently, by the Markov inequality, we have
$\sum_{k=1}^{\infty}P\left(\frac{V_{2^{k}}k}{2^{k}}\geq\epsilon\right)$ $\leq\frac{1}{\epsilon}\sum_{k=1}^{\infty}\frac{\operatorname{\mathbf{E}}V_{2^{k}}k}{2^{k}}$
$\leq\frac{1}{\epsilon}\sum_{k=1}^{\infty}\left(\frac{1}{k^{2}}+kP\left(K_{0}\geq\frac{2^{k}}{k^{3}}\right)\right)<\infty.$ (56)
Hence $\lim_{k\to\infty}\operatorname{\mathbf{E}}V_{2^{k}}k/2^{k}=0$, whereas by the Borel-Cantelli lemma [61, Theorems 4.3], we derive $V_{2^{k}}k/2^{k}<\epsilon$ for all but finitely many $k$ almost surely. Notice now that $V_{t}$ is a growing function of $t$. Hence $(V_{t}\log t)/t\leq 2V_{2^{k}}k/2^{k}$ for $2^{k-1}\leq t\leq 2^{k}$. In particular, we have (52) and $\limsup_{t\to\infty}(V_{t}\log t)/t\leq 2\epsilon$ almost surely. By arbitrariness of $\epsilon$, we derive (53). ∎
4 Information
In this section, we demonstrate propositions that are central from our particular perspective. The first result concerns the flow of information in the description of a sample drawn from a multiperiodic process. We show that a sample of a multiperiodic process carries exactly the same information as the set of seeds for types appearing in the sample. These two objects can be computed from one another and hence their entropies are equal. Consequently, this result allows to sandwich bound the block entropy of a multiperiodic process by the number of types and related statistics. In particular, Hilberg‘s law follows from Heaps‘ law, cf. [4] for a similar result for Santa Fe processes.
4.1 Seed estimator
Let us analyze the relationship between a finite multiperiodic sample and the pool of random seeds in more detail. For bounding the entropy of the sample, it pays off to state this proposition explicitly.
Theorem 13.
Consider a multiperiodic process $(K_{t})_{t\in\mathbb{Z}}$. We claim that:
  1. 1.
    Set $\left\{(k,\Sigma_{k}):k\in\mathcal{V}_{t}\right\}$ is a function of sample $K_{1}^{t}$.
  2. 2.
    Define the seed estimator
    $\hat{\Sigma}_{k}:=\begin{cases}\Sigma_{k},&k\in\mathcal{V}_{t},\\ \pi_{k},&k\not\in\mathcal{V}_{t}.\end{cases}$ (57)
    Let $(\hat{K}_{t})_{t\in\mathbb{Z}}$ be the multiperiodic sequence with seeds $(\hat{\Sigma}_{k})_{k\in\mathbb{N}}$. Then
    $K_{1}^{t}=\hat{K}_{1}^{t}.$ (58)
  3. 3.
    Sample $K_{1}^{t}$ is a function of set $\left\{(k,\Sigma_{k}):k\in\mathcal{V}_{t}\right\}$.
  4. 4.
    If $k\in\mathcal{V}_{t}$ then $k\leq M_{t}$ and $\Sigma_{k}\leq t$, where
    $M_{t}:=\max\mathcal{V}_{t}.$ (59)
Proof.
The subsequent claims are dealt with below.
  • Let $L_{k}=\min\left\{t\geq 1:K_{t}^{(k)}=k\right\}$, where $(K^{(k)}_{t})_{t\in\mathbb{N}}$ is the subsequence of sequence $(K_{t})_{t\in\mathbb{N}}$ from which we have removed all tokens $k_{t}<k$. If $k\in\mathcal{V}_{t}$ then the first $L_{k}$ tokens of $(K^{(k)}_{t})_{t\in\mathbb{N}}$ are a function of sample $K_{1}^{t}$. By claim 2. of Theorem 2, $L_{k}=\Sigma_{k}$. Hence we obtain claim 1.
  • We observe that altering $\Sigma_{k}$ for $k\not\in\mathcal{V}_{t}$ to any value in set
    $\left\{\Sigma_{k},\Sigma_{k}+1,\ldots,\pi_{k}\right\}$
    does not change the resulting multiperiodic sample $K_{1}^{t}$. It is so since any such change only delays the occurrence of token $k$ beyond the scope of $K_{1}^{t}$, leaving the tokens within $K_{1}^{t}$ intact. Hence claim 2. follows.
  • Obviously, claim 2. implies claim 3.
  • If $k\in\mathcal{V}_{t}$ then $k\leq\max\mathcal{V}_{t}$ and $\Sigma_{k}=L_{k}\leq t$ since the first $L_{k}$ tokens of sequence $(K^{(k)}_{t})_{t\in\mathbb{N}}$ are a subsequence of sample $K_{1}^{t}$. Hence we obtain claim 4. ∎
4.2 Block entropy
Now, we will sandwich bound the Shannon entropy in terms of the number of types. For this goal, we will use a bound stemming from the Hölder and Jensen inequalities. Let $\ln$ denote the natural logarithm.
Lemma 1.
Suppose that $X\geq 0$ and $Y\geq 1$. Then for any $p>1$, we have
$\operatorname{\mathbf{E}}X\ln Y\leq(\operatorname{\mathbf{E}}X^{p})^{1/p}\left(\ln\operatorname{\mathbf{E}}Y+\frac{1}{p-1}\right).$ (60)
Proof.
Let $p>1$ and q=p/(p1)q=p/(p-1). For $X,Z\geq 0$, the Hölder inequality is
$\operatorname{\mathbf{E}}XZ\leq(\operatorname{\mathbf{E}}X^{p})^{1/p}(\operatorname{\mathbf{E}}Z^{q})^{1/q}.$ (61)
The second derivative of function $y\mapsto(\ln y)^{q}$ is negative for $y>e^{q-1}$. Hence by the Jensen inequality (J), we obtain
$\operatorname{\mathbf{E}}(\ln Y)^{q}\leq\operatorname{\mathbf{E}}(\ln(e^{q-1}Y))^{q}\stackrel{{\scriptstyle(J)}}{{\leq}}(\ln(e^{q-1}\operatorname{\mathbf{E}}Y))^{q}\leq\left(\ln\operatorname{\mathbf{E}}Y+\frac{1}{p-1}\right)^{q}.$ (62)
Using $Z=\ln Y$ in the Hölder inequality and chaining it with the above chain of inequalities, we derive the claim. ∎
Subsequently, we bound the entropy. Here $\log$ is the base 2 logarithm.
Theorem 14.
For a multiperiodic process $(K_{t})_{t\in\mathbb{Z}}$ and for any $q\geq 1$,
$u_{t}-1\leq H(K_{1}^{t})\leq\left(e\operatorname{\mathbf{E}}V_{t}+3\right)\left((q+2)\log t+q\log\operatorname{\mathbf{E}}K_{0}^{1/q}\right)+2.$ (63)
Proof.
By claim 1. of Theorem 13 and property $k<u_{t}\implies k\in\mathcal{V}_{t}$, the tuple of seeds $\Sigma_{1}^{u_{t}-1}$ is a function of sample $K_{1}^{t}$. Hence we obtain the lower bound
$H(K_{1}^{t})\geq H(\left\{(k,\Sigma_{k}):k\in\mathcal{V}_{t}\right\})\geq H(\Sigma_{1}^{u_{t}-1})\geq u_{t}-1,$ (64)
since seeds are independent and uniformly distributed over two or more values. As for the upper bound, by claims 3. and 4. of Theorem 13, we obtain
$H(K_{1}^{t})$ $\leq H(\left\{(k,\Sigma_{k}):k\in\mathcal{V}_{t}\right\})$
$\leq H(M_{t})+H(\left\{(k,\Sigma_{k}):k\in\mathcal{V}_{t}\right\}|M_{t})$
$\leq 2\operatorname{\mathbf{E}}\log(M_{t}+1)+\operatorname{\mathbf{E}}(V_{t}+1)(\log M_{t}+\log t)$
$\leq 2\operatorname{\mathbf{E}}\log M_{t}+\operatorname{\mathbf{E}}(V_{t}+1)(\log M_{t}+\log t)+2.$ (65)
Next, we observe $1\leq M_{t}^{1/q}\leq\sum_{j=1}^{t}K_{j}^{1/q}$. Hence $\operatorname{\mathbf{E}}M_{t}^{1/q}\leq t\operatorname{\mathbf{E}}K_{0}^{1/q}$ by stationarity. Thus, by Lemma 1, for any $p>1$, we may write
$H(K_{1}^{t})$ $\leq 2\operatorname{\mathbf{E}}q\log M_{t}^{1/q}+\operatorname{\mathbf{E}}(V_{t}+1)\left(q\log M_{t}^{1/q}+\log t\right)+2$
$\leq[(\operatorname{\mathbf{E}}V_{t}^{p})^{1/p}+3]\left(q\log\operatorname{\mathbf{E}}M_{t}^{1/q}+\log t+\frac{\log e}{p-1}\right)+2$
$\leq[(\operatorname{\mathbf{E}}V_{t}^{p})^{1/p}+3]\left((q+1)\log t+q\log\operatorname{\mathbf{E}}K_{0}^{1/q}+\frac{\log e}{p-1}\right)+2.$ (66)
Let us substitute $p=1+1/\ln t$ and apply inequality $V_{t}^{p}\leq t^{p-1}V_{t}$. Then we may rewrite
$H(K_{1}^{t})$ $\leq\left((t^{p-1}\operatorname{\mathbf{E}}V_{t})^{1/p}+3\right)\left((q+1)\log t+q\log\operatorname{\mathbf{E}}K_{0}^{1/q}+\frac{\log e}{p-1}\right)+2$
$=\left((e\operatorname{\mathbf{E}}V_{t})^{\frac{\ln t}{\ln t+1}}+3\right)\left((q+2)\log t+q\log\operatorname{\mathbf{E}}K_{0}^{1/q}\right)+2$
$\leq\left(e\operatorname{\mathbf{E}}V_{t}+3\right)\left((q+2)\log t+q\log\operatorname{\mathbf{E}}K_{0}^{1/q}\right)+2,$ (67)
which is the claimed inequality. ∎
Hence we have a criterion for the vanishing Shannon entropy rate.
Theorem 15.
Suppose that $\operatorname{\mathbf{E}}K_{0}^{1/q}<\infty$ for a $q\geq 1$. Then the multiperiodic process $(K_{t})_{t\in\mathbb{Z}}$ satisfies
$\lim_{t\to\infty}\frac{H(K_{1}^{t})}{t}$ $=0.$ (68)
Proof.
The claim follows by (52) and (63) since $\operatorname{\mathbf{E}}K_{0}^{1/q}<\infty$ for a $q\geq 1$ implies $\operatorname{\mathbf{E}}(\log K_{0})^{3}<\infty$. ∎
5 Examples
Theorem 14 implies that Hilberg‘s law (4) for a multiperiodic process essentially follows from Heaps‘ law (9). In this section, we will work out two concrete examples of processes. First, we will analyze the multiperiodic process with all periods equal to a fixed constant. This case leads to the number of types that grows logarithmically with the sample size. Second, we will inspect the multiperiodic process with periods that grow linearly. This case leads to the number of types that grows like a power law, Heaps‘ law (9) indeed. Consequently, for this example, we also obtain Hilberg‘s law (4).
5.1 Constant periods
First, let us consider the multiperiodic process where periods equal to a fixed constant. To be concrete, let $\pi_{k}=p$ for all $k\in\mathbb{N}$. Then we obtain a geometric distribution of $P(K_{0}=k)=p_{k}$. In particular, we may bound
$f_{k}$ $=\prod_{j=1}^{k-1}\frac{\pi_{j}-1}{\pi_{j}}=\left(\frac{p-1}{p}\right)^{k-1},$ (69)
$p_{k}$ $=f_{k}-f_{k+1}=\frac{1}{\pi_{k}}\prod_{j=1}^{k-1}\frac{\pi_{j}-1}{\pi_{j}}=\frac{1}{p}\left(\frac{p-1}{p}\right)^{k-1},$ (70)
$w_{k}$ $=\sum_{j=2}^{k}\frac{1}{f_{j}}+\frac{\pi_{k}}{f_{k}}=\sum_{j=1}^{k-1}\left(\frac{p}{p-1}\right)^{j}+p\left(\frac{p-1}{p}\right)^{k-1}$
$=\frac{\left(\frac{p}{p-1}\right)^{k}-1}{\frac{p}{p-1}-1}-1+\left(\frac{p}{p-1}\right)^{k-1}p=2p\left(\frac{p}{p-1}\right)^{k-1}-1.$ (71)
Hence we may evaluate
$u_{t}$ $=\min\left\{k:w_{k}>t\right\}=\min\left\{k:2p\left(\frac{p}{p-1}\right)^{k-1}>t+1\right\}$
$=\left\lceil\frac{\log(t+1)-\log p-1}{\log\left(\frac{p}{p-1}\right)}\right\rceil+2,$ (72)
$g_{t}$ $=\max\left\{k:p_{k}^{-1}\leq t\right\}=\max\left\{k:p\left(\frac{p}{p-1}\right)^{k-1}\leq t\right\}$
$=\left\lfloor\frac{\log(t+1)-\log p}{\log\left(\frac{p}{p-1}\right)}\right\rfloor+1,$ (73)
$\log q_{t}$ $=\log f_{g_{t}+1}=-g_{t}\log\left(\frac{p}{p-1}\right)\leq-\log(t+1)+\log p.$ (74)
Hence bound $u_{t}-1\leq\operatorname{\mathbf{E}}V_{t}\leq g_{t}+tq_{t}$ can be rewritten as
$\left\lceil\frac{\log(t+1)-\log p-1}{\log\left(\frac{p}{p-1}\right)}\right\rceil+1\leq\operatorname{\mathbf{E}}V_{t}\leq\left\lfloor\frac{\log(t+1)-\log p}{\log\left(\frac{p}{p-1}\right)}\right\rfloor+1+p.$ (75)
In this case, the Shannon entropy rate vanishes but we do not have Hilberg‘s law (11)—by Theorem 14.
5.2 Linear periods
Now, we consider the multiperiodic process where periods grow approximately in a linear fashion. To be concrete, let $\pi_{k}=1+\left\lceil ck\right\rceil$ for a $c>0$ and all $k\in\mathbb{N}$. We may upper bound
$\prod_{j=1}^{k-1}\frac{\pi_{j}-1}{\pi_{j}}$ $=\prod_{j=1}^{k-1}\left(1-\frac{1}{1+\left\lceil cj\right\rceil}\right)\leq\exp\left(\sum_{j=1}^{k-1}\ln\left(1-\frac{1}{cj+2}\right)\right)$
$\leq\exp\left(-\sum_{j=1}^{k-1}\frac{1}{cj+2}\right)\leq\exp\left(-\int_{1}^{k}\frac{dx}{cx+2}\right)$
$=\exp\left(-\frac{1}{c}\ln\frac{ck+2}{c+2}\right)=\left(\frac{c+2}{ck+2}\right)^{1/c}.$ (76)
Similarly, we may bound conversely
$\prod_{j=1}^{k-1}\frac{\pi_{j}}{\pi_{j}-1}$ $=\prod_{j=1}^{k-1}\left(1+\frac{1}{\left\lceil cj\right\rceil}\right)\leq\exp\left(\sum_{j=1}^{k-1}\ln\left(1+\frac{1}{cj}\right)\right)$
$\leq\exp\left(\sum_{j=1}^{k-1}\frac{1}{cj}\right)\leq\exp\left(\frac{1}{c}+\int_{1}^{k}\frac{dx}{cx}\right)$
$=\exp\left(\frac{1}{c}+\frac{1}{c}\ln k\right)=(ek)^{1/c}.$ (77)
In consequence, we obtain Zipf‘s law (7) for $P(K_{0}=k)=p_{k}$. In particular, we may bound
$f_{k}$ $=\prod_{j=1}^{k-1}\frac{\pi_{j}-1}{\pi_{j}}\geq(ek)^{-1/c},$ (78)
$f_{k}$ $=\prod_{j=1}^{k-1}\frac{\pi_{j}-1}{\pi_{j}}\leq\left(\frac{c+2}{ck+2}\right)^{1/c}\leq\left(\frac{c+2}{c}\right)^{1/c}k^{-1/c},$ (79)
$p_{k}$ $=\frac{1}{\pi_{k}}\prod_{j=1}^{k-1}\frac{\pi_{j}-1}{\pi_{j}}\geq\frac{1}{ck+1}(ek)^{-1/c}\geq\frac{e^{-1/c}}{c+2}k^{-(c+1)/c},$ (80)
$p_{k}$ $=\frac{1}{\pi_{k}}\prod_{j=1}^{k-1}\frac{\pi_{j}-1}{\pi_{j}}\leq\frac{1}{ck+1}\left(\frac{c+2}{ck+2}\right)^{1/c}\leq\frac{1}{c}\left(\frac{c+2}{c}\right)^{1/c}k^{-(c+1)/c}.$ (81)
Thus we may bound the waiting times
$w_{k}$ $=\sum_{j=2}^{k}\frac{1}{f_{j}}+\frac{\pi_{k}}{f_{k}}\leq\sum_{j=2}^{k}(ej)^{1/c}+(ck+2)(ek)^{1/c}$
$\leq\int_{0}^{k}(ex)^{1/c}dx+(ek)^{1/c}+(ck+2)(ek)^{1/c}$
$=\left(\frac{ck}{c+1}+ck+3\right)(ek)^{1/c}\leq\left(c+\frac{c}{c+1}+3\right)e^{1/c}k^{(c+1)/c}.$ (82)
Hence we may bound further
$u_{t}$ $=\min\left\{k:w_{k}>t\right\}\geq\left(\frac{t}{\left(c+\frac{c}{c+1}+3\right)e^{1/c}}\right)^{c/(c+1)},$ (83)
$g_{t}$ $=\max\left\{k:p_{k}^{-1}\leq t\right\}\leq\left(\frac{t}{c}\left(\frac{c+2}{c}\right)^{1/c}\right)^{c/(c+1)},$ (84)
$g_{t}$ $=\max\left\{k:p_{k}^{-1}\leq t\right\}\geq\left(\frac{te^{-1/c}}{c+2}\right)^{c/(c+1)}$ (85)
so as to derive
$q_{t}$ $=f_{g_{t}+1}\leq\left(\frac{c+2}{c}\right)^{1/c}\left(\frac{te^{-1/c}}{c+2}\right)^{-1/(c+1)}.$ (86)
As a result, bound $u_{t}-1\leq\operatorname{\mathbf{E}}V_{t}\leq g_{t}+tq_{t}$ implies Heaps‘ law
$C_{1}t^{c/(c+1)}\leq\operatorname{\mathbf{E}}V_{t}\leq C_{2}t^{c/(c+1)}$ (87)
for certain constants $C_{1},C_{2}>0$. By Theorem 14, to derive Hilberg‘s law (11) with $\beta=c/(c+1)$, it suffices to show that $\operatorname{\mathbf{E}}K_{t}^{1/q}<\infty$ for some $q\geq 1$. In fact, for any $q>c$, we obtain
$\operatorname{\mathbf{E}}K_{t}^{1/q}$ $=\int_{0}^{\infty}P(K_{t}^{1/q}\geq k)dk=\int_{0}^{\infty}P(K_{t}\geq k^{q})dk$
$=\int_{0}^{\infty}f_{\left\lceil k^{q}\right\rceil}dk\leq\int_{1}^{\infty}\left(\frac{c+2}{c}\right)^{1/c}k^{-q/c}dk<\infty.$ (88)
Figure 1: The number of types $V_{t}$ as a function of sequence length $t$ for periods $\pi_{k}=1+\left\lceil ck\right\rceil$ and seeds $\sigma_{k}=1$. The lines are theoretical predictions $V_{t}\sim t^{c/(c+1)}$, where the proportionality constant is chosen by least squares.
To provide some visualization of these results, in Figure 1, we present the growth of the actual number of types $V_{t}$ for three multiperiodic sequences with periods $\pi_{k}=1+\left\lceil ck\right\rceil$ and seeds $\sigma_{k}=1$, where $c\in\left\{0.5,1.0,2.0\right\}$. We notice a good agreement with the theoretical prediction $\operatorname{\mathbf{E}}V_{t}\sim t^{c/(c+1)}$.
6 Conclusion
The example of multiperiodic processes extends mathematical analysis of Hilberg‘s law—the power law growth of entropy for certain stationary sources, possibly entailing statistical language models [3, 11]. This topic is relevant not only for quantitative linguistics [2, 34] but also for large language models in machine learning, via the neural scaling law [5, 6, 7, 8, 4]. As it was shown in the previous research revolving around the example of Santa Fe processes [2, 34, 35, 43], the topic can be linked to more traditional analyses of Zipf‘s law [36, 37]. This study has supplied further examples and findings. We hope that the earlier constructed Santa Fe processes and the multiperiodic processes described here may be applied as toy data sources to be compared with natural data in the study of language and artificial intelligence.
Prior versions
This work stems from rethinking and heavily editing an imperfect manuscript [67], which was posted to ArXiv on February 17, 2023. That manuscript has not been posted or published anywhere else but was presented as a seminar talk at the Math Machine Learning seminar in the Max Planck Institute for Mathematics in the Sciences.
Funding
This work received no external funding.
Acknowledgment
I thank Jan Mielniczuk, Szymon Jaroszewicz, Marcus Hutter, and Guido Montúfar for discussing prior versions of this paper.